### v1

# Improving Spiking Neural Network Accuracy With Color Model Information Encoded Bit Planes

Abstract
Spiking neural networks (SNNs) have emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint. However, their practical adoption is constrained by several challenges, prominently among them being performance optimization. In this study, we present a novel approach to enhance the performance of SNNs through a new encoding method that exploits bit planes derived from various color models of input image data for spike encoding. Our proposed technique is designed to improve the computational accuracy of SNNs compared to conventional methods without increasing model size. Through extensive experimental validation, we demonstrate the effectiveness of our encoding strategy in achieving performance gain across multiple computer vision tasks. To the best of our knowledge, this is the first research endeavor applying color spaces within the context of SNNs. By leveraging the unique characteristics of color spaces, we hope to unlock new potentials in SNNs performance, potentially paving the way for more efficient and effective SNNs models in future researches and applications.
Index Terms:
Spiking neural network, spike encoding, image classification
I Introduction
The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[1], largely driven by the development and optimization of deep learning algorithms. Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[2]. To bridge this gap, spiking neural networks (SNNs) have emerged as a promising alternative, leveraging the temporal dimension of neuronal activity to more closely emulate the behavior of biological brains.
SNNs utilize discrete spike events to transmit information between neurons, mirroring the asynchronous and event-driven communication observed in biological neurons[3]. This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[4]. The inherent sparsity and temporal coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [5].
Despite their potential, SNNs pose significant challenges, particularly in training methodologies and network optimization. Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[6]. Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9], have been developed to address these challenges, enabling the effective training of SNNs.
Prior research in ANNs has extensively explored the application of image color spaces to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing[10, 11]. Motivated by these advancements, we aim to investigate whether these insights on color spaces can similarly improve the performance of SNNs. This led us to formulate our research questions:
  • Does encoding color spaces encapsulate the necessary information to enhance the performance of SNNs?
  • How can we encode color spaces information into SNNs?
To answer these hypotheses, we observed that the bit planes of images contain pulse-like information, which has been utilized previously on ANNs [12] and can be apply to encode color spaces information for SNNs to possibly improve accuracy.
II Spiking Neural Network
In this section we will cover some literature survey related to our research, particularly about SNNs in general and some common knowledge on SNN architecture utilized in our experiments.
II-A History of SNNs
Phenomenological neuron models can be progressively simplified until reaching a stage where even neural spikes are not modeled, leading to first and second generation neurons [13, 14]. This simplification significantly reduces computational demands but also decreases biological realism. The detail level in modeling often correlates with available computing power. For example, ANNs based on second generation neurons were historically limited by computing power [15], restricting the size of networks and the complexity of neuronal models. This limitation influenced the design of learning algorithms. Even with advancements in computing power, the complexity of neuronal models did not proportionately increase because existing learning algorithms were not compatible with more detailed models. Consequently, two distinct research fields emerged.
  • ANNs[16] focused on large networks of simplified neuron-like units, capable of learning through algorithms like backpropagation, and were successfully applied in pattern recognition, classification, and pattern completion tasks.
  • Computational Neuroscience[17] in contrast used detailed biophysical and phenomenological models in smaller networks to study electrophysiological processes, pattern generation, and dynamic behaviors of small neuron groups.
Although there have been attempts to create large networks of detailed biophysical neuron models [18], it remains challenging due to computational constraints.
Recent advances in computing power have blurred the lines between these fields. ANNs have become more biologically realistic, while networks of biophysical neurons have grown larger and more detailed. However, computational power still limits the application of detailed models in large neural networks for pattern recognition and classification. With improved computing power, new learning algorithms are being developed for these models. SNNs, which use spiking neurons modeled by phenomenological models like the Spike Response Model[19], represent the next step towards this goal. Due to the computational burden, biophysical models are less common in certain SNN applications. This paper focuses on research involving SNNs with phenomenological models, excluding those using biophysical models.
II-B The Integrate-and-Fire model
One of the earliest neuron models is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [20]. In this model, the neuron is characterized by its membrane potential $V$:
$C\frac{dV(t)}{dt}=I(t)$ (1)
which evolves over time in response to an input current $I(t)$. This evolution is described by the time derivative of the capacitance law, $Q=CV$, where $C$ is the membrane capacitance. As an input current is applied, the membrane potential increases until it reaches a fixed threshold $V_{\text{th}}$. At this point, the neuron generates a spike modeled as a delta function, and the membrane potential is reset to its resting value, after which the process repeats. In this model, the firing rate increases linearly and indefinitely with the increase in input current.
While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons, challenges still remain in the conversion process. One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches. Despite the introduction of both data-driven and model-driven threshold determination methods[21], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts. Additionally, IF neurons lack time-dependent memory. Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires. This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [22].
II-C Deep Residual Learning in Spiking Neural Networks
The success of ResNet[23] in deep learning has motivated efforts to extend residual learning to SNN. Prior attempts, commonly referred to as Spiking ResNets[24], have largely replicated the architecture of ResNet[23] by replacing ReLU activation functions with spiking neurons. However, this approach presents certain limitations:
  • Inability to achieve identity mapping across neuron models: Residual learning is predicated on the concept of identity mapping, where layers can be trained to approximate an identity function. While this is straightforward in ResNets[23] with ReLU activation, it is not always feasible in Spiking ResNets[24]. Specifically, [25] claimed that for certain spiking neuron models, such as leaky integrate-and-fire (LIF) neurons[20] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging. This complicates the realization of identity mapping.
  • Vanishing/exploding gradients: Even in cases where identity mapping can be achieved, Spiking ResNets[24] are prone to vanishing or exploding gradients during training. This occurs because the gradient propagation through multiple layers of spiking neurons may either diminish to zero or grow uncontrollably, complicating the training of deeper networks.
To address these issues, SEW ResNet[25] was introduced as a replacement design for residual blocks while utilizing IF neurons. The SEW block capitalizes on the binary nature of spikes by applying element-wise functions $g$ to the output of the residual mapping $A_{l}[t]$ and the input spike train $S_{l}[t]$[25]. This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models. With these advantages in mind, we have decide to choose (but not limited to) SEW Resnet[25] architecture as our main experimenting architecture.
III Spike Coding in SNNs
In the context of SNN there are several encoding methods, each with unique advantages and disadvantages.
  • Rate coding [26, 27] is the predominant coding technique utilized for spikes in SNNs. In rate coding, the number of spikes within a specific time interval is counted, and the firing rate is interpreted as the signal intensity. This method inherently includes a high level of redundancy and exhibits tolerance to a certain degree of noise.
  • Temporal coding[3] is the form of coding where spike patterns are distributed over time. Population-based coding methods [28] encode information based on the number of spikes within a given time frame. In contrast, inter-spike interval methods [29] use the time between two spikes as their coding scheme. Temporal coding allows for arbitrarily sized data patterns, though longer patterns require extended time periods.
We primarily focus on rate coding, as it is one of the most widely used encoding schemes in conventional SNNs for representing signals with spike trains. In the rate coding scheme, the firing frequency of a neuron represents the signal intensity. The firing rate $r_{i}$ of the i-th neuron provided by [30] can be expressed as follows:
$r_{i}=\frac{N_{s}}{T}=\frac{\sum_{t=1}^{T}s_{i}(t)}{T}$ (2)
When $N_{s}$ spikes are fired over T time steps, each time step can either produce a spike or none. The firing rate is 1 if a neuron fires at every time step, and 0 if it does not fire at all. In image classification, for instance, if a pixel intensity ranges [0.0, 1.0], it can be represented by a spike train with a corresponding firing rate [0.0, 1.0] as input to the SNN.
In this research, we will be using the Poisson-distributed rate coding [30, 31] scheme where spike $s_{i}$ for i-th pixel with intensity $P_{i}$ can be interpreted as:
$s_{i}=\begin{cases}1&\text{if $random(0.0,1.0)\geq P_{i}$}\\ 0&\text{otherwise}\end{cases}$ (3)
IV Color spaces and their application in deep learning
This section will provide an overview of relevant concepts for our research, focusing specifically on the color models explored and their applications in deep learning algorithms.
IV-A Color spaces
Color spaces are fundamental in computer graphics as they define how colors are represented, manipulated, and displayed within digital images. In computer graphics applications, understanding and effectively managing color spaces is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces. In this subsection, we will discuss the color spaces under investigation in our experiments, divided by certain criteria from [32, 33], including: device-oriented models, user-oriented models and device-independent models.
IV-A1 Device-oriented models
These models are influenced by the signals of the devices and the resulting colors are affected by the tools used for display. Widely utilized in various applications that require color consistency with hardware tools, examples include any hardware devices used for human visual perception, such as televisions and video systems . While there are different models for separate use-cases and devices, for now we will only concern with RGB, CMY(K) and YCbCr.
IV-A2 User-oriented models
User-oriented color models are considered as a pathway between the observer and the device handling the color information. These models enable users to describe and approximate their color perceptions from the presented colors. For our experiments, we mostly investigate cylindrically described models, including HSL and HSV. According to [32], these models also inherit other special properties such as devices dependence and cylindrical-coordinate representation.
IV-A3 Device-independent models
Device-independent color models are essential for defining color signals without reliance on the specific characteristics of any particular device or application. These models play a crucial role in applications that involve color comparisons and the transmission of visual information across networks that connect various hardware platforms. We will mostly research color spaces defined by the International Commission on Illumination (CIE), including CIE XYZ and CIE LAB.
IV-B Usage of color models in deep learning
Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks. For example, research on traffic light recognition[34] tested a range of color models across different deep learning architectures, with results indicating that the RGB model remains favored despite some of its limitations. This suggests that, while RGB is not perfect, it still outperforms other models in certain applications.
In another study focused on applying convolutional neural networks (CNNs) for skin detection against adversarial attacks of generative models[35], RGB and YCbCr color models demonstrated superior performance, particularly in handling varying lighting conditions. This highlights the adaptability of these color models in scenarios where lighting inconsistencies can degrade model accuracy.
Moreover, the development of ColorNet[10] illustrates the potential of leveraging multiple color spaces to capture a broader set of features. By combining information from these diverse color spaces, ColorNet outperforms traditional models that rely on a single color space. However, this approach introduces computational overhead due to the inclusion of smaller submodels, which can be a challenge despite the model’s increased efficiency in some cases.
Despite extensive researches on the use of color models in ANNs, their application in SNNs remains underexplored and requires further investigation.
V Bit planes application in deep learning
Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance. In one study, researchers used bit planes to encode the input layer in Binary Neural Networks (BNNs)[12]. They argued that bit planes retain crucial spatial information, enabling a reduction in the number of multiply-accumulate operations (MACs) without compromising the model’s accuracy or robustness. This demonstrates the potential of bit planes in streamlining computations while maintaining competitive performance.
Another study explored the use of bit plane slicing on RGB images for adversarial robustness[36]. By selecting the most significant bit planes, the researchers developed a classification model that was better equipped to handle adversarial attacks. This approach capitalizes on the spatial information encoded in different bit planes, which helps the model resist manipulations that would otherwise degrade its performance.
In the domain of medical imaging, bit plane slicing has also been applied to improve accuracy in breast cancer recognition tasks[37]. Researchers found that lower-order bit planes are more susceptible to noise, and by removing them before inputting the data into a convolutional neural network (CNN), they achieved higher classification accuracy. This method highlights how bit plane analysis can be tailored to specific challenges, such as noise reduction in medical images, to enhance model effectiveness.
Similar to color models, although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still relatively unexplored and demands further research.
VI Experiments setup
In this section, we will discuss in details about the implementation of our encoding algorithm and setup for experimenting with this method.
VI-A Preliminary testing
As theoretically mentioned in [38] and practically measured in [39, 40], a Poisson-distributed time steps appear to positively correlate with training accuracy in SNNs and to achieve a correct spike distribution, the input must be sampled to a certain frequency. However in the case of a high time steps input, this approach sacrifices the energy efficiency of SNNs, which is one of their primary advantages. Despite this trade-off, the improvement in training accuracy suggests that the benefits in performance might outweigh the costs in specific applications where accuracy is paramount.
We previously hypothesize that the information contained in the bit planes of color models can be effectively used to encode inputs for SNN training. Bit planes decompose an image into binary levels, each representing a specific bit of the pixel values. This decomposition can potentially preserve critical information from the original image while providing a format that might be more suitable for the spike-based processing in SNNs. The granularity of information available in bit planes offers a unique approach to encoding visual data, which could enhance the learning capabilities of SNNs by providing a richer, multi-level representation of the input.
To test our hypotheses, we conducted preliminary experiments using the CIFAR-10 dataset, employing the bit plane encoding method. We compared this approach with traditional rate coding and a combination of both methods (illustrated in details in Figure 2 and Section VI-B) with a rate coding step size $T=10$ (illustrated in Figure 1 and Table I).
Figure 1: Compare training progression on CIFAR-10 [41] between using traditional rate coding (denoted blue), encoding using only bit planes information (denoted orange) and combining bit planes information with rate coding (denoted green).
Encoding method Validation Accuracy (%)
Poisson rate coding with time step T=10 70.69
RGB bit plane encoding 73.49
Poisson rate coding with time step T=10 and combined RGB bit plane encoding 37.28
TABLE I: Validation accuracy comparison on CIFAR-10 [41] between traditional rate coding, encoding using only bit planes information and combining bit planes information with rate coding.
Visible from our results, feeding bit plane information of the encoded RGB color model to the SNNs allowed the model to converge to a certain level of accuracy. Still, the accuracy of purely bit planes encoded inputs still fall short compare to rate coding. On the other hand, a combination of rate coding and bit planes encoding experience an increase in performance, this outcome demonstrates that bit planes in color models contain valuable information that SNNs can learn from to improve their classification ability and encourage us to perform testing on other datasets with different color models (as mentioned in Section IV-A).
VI-B Overall encoding scheme with color spaces
Figure 2: Our proposed encoding method (lower) compare with traditional rate coding method (upper).
As we can see in Figure 2, for our research the encoding scheme adopted integrates a dual approach featuring both traditional rate coding and bit plane encoding methodologies. Initially, the process commences with the conventional rate coding procedure applied to the image tensor. Subsequently, depending on the chosen color model (for example RGB), the image tensor may undergo conversion to the specified model or proceed directly to encoding if already in RGB format. Following this, the bit plane encoding algorithm (will be discussed in details in Section VI-C and Algorithm 1) is employed to transform the color-modeled image into binary planes. The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from rate coding and bit plane encoding. These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes, facilitating enhanced performance in our computational models.
VI-C Bit plane encoding algorithm
To achieve our research objectives, we have developed Algorithm 1 designed to encode an image tensor batch $X$ into $n_{bit}$ bit planes. To determine $n_{bit}$, it is essential to know the maximum possible value $X_{max}$ of the tensor batch. Since different color models have varying $X_{max}$, this value must be predefined according to the selected color model (for example, if RGB is chosen, then $X_{max}=255$). Once $X_{max}$ is defined, $n_{bit}$ can be calculate as $\left\lceil\log_{2}(X_{max})\right\rceil$.
To calculate the bit planes of an unnormalized integer input tensor batch $X$, we perform element-wise division of $X$ by 2. The remainder from this division constitutes a bit plane of $X$, while the quotient is saved for the extraction of subsequent bit planes. This process is repeated until we obtain $n_{bit}$ bit planes. The resulting bit planes are then concatenated to form a binary tensor batch with the shape $(n_{bit},B,C,H,W)$. This format closely resembles that of rate encoding and ensures determinism in the encoding results without having to sample more than $n_{bit}$ times.
Algorithm 1 Bit plane encoding algorithm
$X_{shape}\in(B,C,H,W)$triangleright Constraining input format of multiple images. If input is a single image we should have $B=1$.
$X_{max}$ is defined
function BiplaneEncode($X,X_{max}$)
     $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$.
     $L\leftarrow[\text{ }]$triangleright Empty list initialization.
     for $i\leftarrow 0,i<n_{bit},i++$ do
         $L.insert(X\mod{2})$triangleright Element-wise modulo of tensor, then store the bit plane.
         $X\leftarrow\left\lfloor X/2\right\rfloor$triangleright Element-wise division of tensor.
     end for
     return $concat(L)$triangleright Return a concatenated tensor batches of new dimension $(n_{bit},B,C,H,W)$
end function
VI-D Experiment setup
To investigate performance gains during training process, the experimented model should be a medium-sized SNN with potential for improvement in computer vision tasks. After careful consideration, we had decided to select SEW-ResNet18 with ADD element-wise function [25] which is a novel medium size SNN tested on various computer vision tasks and suitable to our specifications. The model and experiment process was implemented using PyTorch [42] and SpikingJelly[43], incorporating color conversion algorithms based on implementations from Kornia [44]. During the rate coding steps, all training processes utilized step size $T=10$. For optimization, we employed the Adam optimizer [45] with a learning rate of $lr=1e-3$ and $\beta=(0.9,0.999)$ for all tested datasets. Cross-Entropy Loss with Softmax function and mean reduction $L(x,y)$[42] across $N$ mini-batch dimension was also widely employed as the loss function:
$L(x,y)=\frac{\sum_{n=1}^{N}l_{n}}{N},$ (4)
$l_{n}=-\sum_{c=1}^{C}w_{c}\log\frac{\exp{x_{n,c}}}{\sum_{i=1}^{C}\exp{x_{n,i}}}y_{n,c}$
with $l_{n}$ is loss of each mini-batch, $x$ is the input, $y$ is the target, $w$ is the weight, total amount of class $C$ where mini-class $c,i\in C$.
VII Results
In this section we will discuss results of our main experiments along with some interesting findings beside original hypotheses.
VII-A The effect of bit plane encoding on colored images
For our experiment, we conducted tests on various image classification datasets, including CIFAR10[41], CIFAR100[41], EuroSAT[46], Caltech101[47], Caltech256[48], Imagenette[49] (which is a subset of ImageNet[50]), GTSRB[51], and Food101[52]. We allocated 80% of each dataset for training and the remaining 20% for validation. All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU, where CIFAR10[41], CIFAR100[41] and EuroSAT[46] are trained on original input size while input of all other dataset are resize to 224x224 pixels.
As depicted in Table II, the average accuracy and loss curves across all tested color models demonstrated improved performance. Despite a deterministic increase in inference time and input size, the results indicate that the information extracted from color model bit planes enhanced the training process. This enhancement serves as an information supplement, leading to higher performance overall without increasing the model size.
Accuracy of encoding methods (%)
Dataset Poisson rate coding With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
CIFAR10[41] 70.69 73.49 73.06 73.64 73.39 73.06 73.63 74.06
CIFAR100[41] 38.57 42.15 41.31 41.62 41.94 41.91 41.73 42.03
Caltech101[47] 61.67 64.55 65.65 65.53 66.46 68.41 65.30 66.22
Caltech256[48] 30.04 41.51 40.65 40.84 40.55 39.94 40.50 40.79
EuroSAT[46] 84.39 88.48 86.61 88.11 88.74 88.81 88.87 87.48
Imagenette[49] 72.99 78.68 77.35 78.78 78.55 78.50 77.35 78.70
Food101[52] 19.62 36.95 37.71 36.98 38.03 36.41 37.10 35.87
GTSRB[51] 93.36 95.35 51.36 70.68 79.18 80.86 51.78 51.78
TABLE II: Comparison of best validation accuracy in percentage between traditional rate coding method and our encoding methods, best performing model is emphasized.
VII-B The effect of bit plane encoding on low resolution grayscale images
Accuracy of encoding methods (%)
Dataset Poisson rate coding with time step T=10 Poisson rate coding with time step T=10 and combined RGB bit plane encoding RGB bit plane encoding with no rate coding
MNIST[53] 98.43 98.51 97.63
Fashion-MNIST[54] 87.99 88.78 85.76
KMNIST[55] 93.85 94.60 91.39
TABLE III: Comparison of best validation accuracy in percentage between encoding methods, best performing model is emphasized.
Our encoding scheme is designed to leverage color model information to enhance the performance of SNN models, which is why testing on grayscale images was not initially prioritized. However, given that bit-plane encoding can be feasibly applied to single-channel grayscale images, we aim to investigate whether this approach yields any significant or insightful results. Although grayscale images consist of a pixel intensity range between [0, 255], we can adapt our encoding process by applying RGB-based encoding with a reduced number of channels. This method would produce eight bit-plane channels that represent the image information in binary form.
As demonstrated in Table III (for further details see Figure 5 in Appendix References), the combination of Poisson rate coding with bit-plane encoding outperformed other coding methods in terms of validation accuracy. Notably, using only bit-plane information resulted in only a slight reduction in accuracy. As discussed in the previous section, we hypothesize that although bit-plane encoding contains less information compared to traditional rate coding, the model can still converge to a certain extent in computer vision tasks. We further suspect that this performance deficit may be less pronounced in less complex tasks (in this situation, is a low resolution grayscale image recognition task).
VIII Conclusion
In conclusion, our research introduces a novel encoding method for SNN that leverages bit planes derived from various color models of input image data. This approach prove to enhances the computational accuracy of SNN without increasing the model size. Through extensive experimental validation, we have demonstrated the effectiveness of our encoding strategy across multiple computer vision tasks, achieving notable performance gains compared to conventional methods. This study is pioneering in its application of color spaces within the context of spiking neural networks, and we believe that our findings open new avenues for the development of more efficient and effective SNN models. Our work sets a foundation for future research and applications, potentially advancing the field of computational neuroscience and artificial intelligence.

### v2

@@
-Improving Spiking Neural Network Accuracy With Color Model Information Encoded Bit Planes
+Improvement of Spiking Neural Network with Bit Planes and Color Models
 Abstract
-Spiking neural networks (SNNs) have emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint.
+Spiking neural network (SNN) has emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint.
 However, their practical adoption is constrained by several challenges, prominently among them being performance optimization.
-In this study, we present a novel approach to enhance the performance of SNNs through a new encoding method that exploits bit planes derived from various color models of input image data for spike encoding.
-Our proposed technique is designed to improve the computational accuracy of SNNs compared to conventional methods without increasing model size.
-Through extensive experimental validation, we demonstrate the effectiveness of our encoding strategy in achieving performance gain across multiple computer vision tasks.
-To the best of our knowledge, this is the first research endeavor applying color spaces within the context of SNNs.
-By leveraging the unique characteristics of color spaces, we hope to unlock new potentials in SNNs performance, potentially paving the way for more efficient and effective SNNs models in future researches and applications.
+In this study, we present a novel approach to enhance the performance of SNN for images through a new coding method that exploits bit plane representation.
+Our proposed technique is designed to improve the accuracy of SNN without increasing model size.
+Also, we investigate the impacts of color models of the proposed coding process.
+Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets.
+To the best of our knowledge, this is the first research that considers bit planes and color models in the context of SNN.
+By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more efficient and effective SNNs models in future researches and applications.
 Index Terms:
-Spiking neural network, spike encoding, image classification
+Spiking neural network, spike coding, image classification
 I Introduction
-The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[1], largely driven by the development and optimization of deep learning algorithms.
+The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[1], largely driven by the developments of deep learning algorithms.
 Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[2].
@@
 Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9], have been developed to address these challenges, enabling the effective training of SNNs.
-Prior research in ANNs has extensively explored the application of image color spaces to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing[10, 11].
-Motivated by these advancements, we aim to investigate whether these insights on color spaces can similarly improve the performance of SNNs.
+Prior research in ANN has explored the application of bit planes (e.g.
+[10]) and color models [11, 12] to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing.
+Motivated by these advancements, we aim to investigate whether bit planes and color models can similarly improve the performance of SNNs.
 This led us to formulate our research questions:
-  • Does encoding color spaces encapsulate the necessary information to enhance the performance of SNNs?
-  • How can we encode color spaces information into SNNs?
-To answer these hypotheses, we observed that the bit planes of images contain pulse-like information, which has been utilized previously on ANNs [12] and can be apply to encode color spaces information for SNNs to possibly improve accuracy.
+  • Does bit plane representation enhance the performance of SNN?
+  • What are impacts of different color models on SNN?
+Especially, we observe that bit planes of an image contain pulse-like information, and can be applied to encode color information for SNNs to possibly improve accuracy.// After extensive experimentation, we concluded that bit planes contain valuable information that can be utilized in the rate coding process as a supplementary source to enhance the accuracy of direct surrogate backpropagation process [13] in SNN architectures.
+Furthermore, bit plane information proved to be reusable across multiple color models in most of the cases examined in our study.
+A preliminary version of this paper is published in [14], where the benefit of including bit plane coding is proved by three grayscale datasets.
+In this extended version, the following new points have been added.
+First, the proposed method was updated to support color images.
+Second, eight more datasets are included.
+Third, impacts of color models are investigated.
 II Spiking Neural Network
-In this section we will cover some literature survey related to our research, particularly about SNNs in general and some common knowledge on SNN architecture utilized in our experiments.
+In this section we present an overview of SNN knowledge related to our research.
 II-A History of SNNs
@@
 This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [24].
+Still, within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better LIF neurons[24], with the SEW-ResNet[25] architecture being one such example.
+Figure 1: Illustrated bit planes of a sample image from CIFAR10[26] dataset.
 II-C Deep Residual Learning in Spiking Neural Networks
@@
 This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models.
-With these advantages in mind, we have decide to choose (but not limited to) SEW Resnet[25] architecture as our main experimenting architecture.
-III Spike Coding in SNNs
-In the context of SNN there are several encoding methods, each with unique advantages and disadvantages.
-  • Rate coding [26, 27] is the predominant coding technique utilized for spikes in SNNs.
-In rate coding, the number of spikes within a specific time interval is counted, and the firing rate is interpreted as the signal intensity.
-This method inherently includes a high level of redundancy and exhibits tolerance to a certain degree of noise.
-  • Temporal coding[3] is the form of coding where spike patterns are distributed over time.
-Population-based coding methods [28] encode information based on the number of spikes within a given time frame.
-In contrast, inter-spike interval methods [29] use the time between two spikes as their coding scheme.
-Temporal coding allows for arbitrarily sized data patterns, though longer patterns require extended time periods.
-We primarily focus on rate coding, as it is one of the most widely used encoding schemes in conventional SNNs for representing signals with spike trains.
-In the rate coding scheme, the firing frequency of a neuron represents the signal intensity.
-The firing rate $r_{i}$ of the i-th neuron provided by [30] can be expressed as follows:
-$r_{i}=\frac{N_{s}}{T}=\frac{\sum_{t=1}^{T}s_{i}(t)}{T}$ (2)
-When $N_{s}$ spikes are fired over T time steps, each time step can either produce a spike or none.
-The firing rate is 1 if a neuron fires at every time step, and 0 if it does not fire at all.
-In image classification, for instance, if a pixel intensity ranges [0.0, 1.0], it can be represented by a spike train with a corresponding firing rate [0.0, 1.0] as input to the SNN.
-In this research, we will be using the Poisson-distributed rate coding [30, 31] scheme where spike $s_{i}$ for i-th pixel with intensity $P_{i}$ can be interpreted as:
-$s_{i}=\begin{cases}1&\text{if $random(0.0,1.0)\geq P_{i}$}\\ 0&\text{otherwise}\end{cases}$ (3)
-IV Color spaces and their application in deep learning
-This section will provide an overview of relevant concepts for our research, focusing specifically on the color models explored and their applications in deep learning algorithms.
-IV-A Color spaces
-Color spaces are fundamental in computer graphics as they define how colors are represented, manipulated, and displayed within digital images.
-In computer graphics applications, understanding and effectively managing color spaces is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces.
-In this subsection, we will discuss the color spaces under investigation in our experiments, divided by certain criteria from [32, 33], including: device-oriented models, user-oriented models and device-independent models.
-IV-A1 Device-oriented models
-These models are influenced by the signals of the devices and the resulting colors are affected by the tools used for display.
-Widely utilized in various applications that require color consistency with hardware tools, examples include any hardware devices used for human visual perception, such as televisions and video systems .
-While there are different models for separate use-cases and devices, for now we will only concern with RGB, CMY(K) and YCbCr.
-IV-A2 User-oriented models
-User-oriented color models are considered as a pathway between the observer and the device handling the color information.
-These models enable users to describe and approximate their color perceptions from the presented colors.
-For our experiments, we mostly investigate cylindrically described models, including HSL and HSV.
-According to [32], these models also inherit other special properties such as devices dependence and cylindrical-coordinate representation.
-IV-A3 Device-independent models
-Device-independent color models are essential for defining color signals without reliance on the specific characteristics of any particular device or application.
-These models play a crucial role in applications that involve color comparisons and the transmission of visual information across networks that connect various hardware platforms.
-We will mostly research color spaces defined by the International Commission on Illumination (CIE), including CIE XYZ and CIE LAB.
-IV-B Usage of color models in deep learning
-Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks.
-For example, research on traffic light recognition[34] tested a range of color models across different deep learning architectures, with results indicating that the RGB model remains favored despite some of its limitations.
-This suggests that, while RGB is not perfect, it still outperforms other models in certain applications.
-In another study focused on applying convolutional neural networks (CNNs) for skin detection against adversarial attacks of generative models[35], RGB and YCbCr color models demonstrated superior performance, particularly in handling varying lighting conditions.
-This highlights the adaptability of these color models in scenarios where lighting inconsistencies can degrade model accuracy.
-Moreover, the development of ColorNet[10] illustrates the potential of leveraging multiple color spaces to capture a broader set of features.
-By combining information from these diverse color spaces, ColorNet outperforms traditional models that rely on a single color space.
-However, this approach introduces computational overhead due to the inclusion of smaller submodels, which can be a challenge despite the model’s increased efficiency in some cases.
-Despite extensive researches on the use of color models in ANNs, their application in SNNs remains underexplored and requires further investigation.
-V Bit planes application in deep learning
+SEW ResNet [25] was also the first work to propose and apply a surrogate gradient approach for optimizing SNNs.
+Using backpropagation methods to train SNNs [29, 13] has become a common practice in neuromorphic computing.
+However, applying backpropagation in SNNs presents certain challenges.
+In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
+$H(\theta,t)=\begin{cases}1&\text{if $\theta-t\geq 0$}\\ 0&\text{otherwise}\end{cases}$ (2)
+The derivative $\partial H$ of this function is given by:
+$\frac{\partial H}{\partial\theta}=\begin{cases}+\infty&\text{if $\theta-t=0$}\\ 0&\text{otherwise}\end{cases}$ (3)
+As shown, directly differentiating $H(\theta,t)$ results in unstable gradients.
+To address this, [25] proposed substituting the derivative of the Heaviside function $\partial H$ with a numerically stable gradient from a different activation function during backpropagation, such as the derivative $\partial\sigma$ of the Sigmoid function $\sigma(\theta,t)$:
+$\sigma(\theta,t)=\frac{1}{1+e^{\theta-t}}$ (4)
+$\frac{\partial H}{\partial\theta}=\frac{\partial\sigma}{\partial\theta}=\frac{e^{\theta-t}}{(e^{\theta-t}+1)^{2}}$
+This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively.
+With these advantages, we decided to choose SEW Resnet[25] architecture for our study.
+III Bit Planes and Color Models in Deep Learning
+This section presents an overview of bit planes and color models, along with their applications in deep learning algorithms.
+III-A Bit planes in deep learning
 Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance.
-In one study, researchers used bit planes to encode the input layer in Binary Neural Networks (BNNs)[12].
+In [10], researchers used bit planes to encode the input layer in Binary Neural Network (BNN).
 They argued that bit planes retain crucial spatial information, enabling a reduction in the number of multiply-accumulate operations (MACs) without compromising the model’s accuracy or robustness.
 This demonstrates the potential of bit planes in streamlining computations while maintaining competitive performance.
-Another study explored the use of bit plane slicing on RGB images for adversarial robustness[36].
-By selecting the most significant bit planes, the researchers developed a classification model that was better equipped to handle adversarial attacks.
+The use of bit plane slicing on RGB images for adversarial robustness was explored in [30].
+By selecting the most significant bit planes, the authors developed a classification model that was better equipped to handle adversarial attacks.
 This approach capitalizes on the spatial information encoded in different bit planes, which helps the model resist manipulations that would otherwise degrade its performance.
@@
 This method highlights how bit plane analysis can be tailored to specific challenges, such as noise reduction in medical images, to enhance model effectiveness.
-Similar to color models, although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still relatively unexplored and demands further research.
-VI Experiments setup
-In this section, we will discuss in details about the implementation of our encoding algorithm and setup for experimenting with this method.
-VI-A Preliminary testing
-As theoretically mentioned in [38] and practically measured in [39, 40], a Poisson-distributed time steps appear to positively correlate with training accuracy in SNNs and to achieve a correct spike distribution, the input must be sampled to a certain frequency.
-However in the case of a high time steps input, this approach sacrifices the energy efficiency of SNNs, which is one of their primary advantages.
-Despite this trade-off, the improvement in training accuracy suggests that the benefits in performance might outweigh the costs in specific applications where accuracy is paramount.
-We previously hypothesize that the information contained in the bit planes of color models can be effectively used to encode inputs for SNN training.
-Bit planes decompose an image into binary levels, each representing a specific bit of the pixel values.
-This decomposition can potentially preserve critical information from the original image while providing a format that might be more suitable for the spike-based processing in SNNs.
+Although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still unexplored.
+As far as we concern, our research is first to explore application of bit planes on SNNs spike coding process.
+III-B Color models in deep learning
+Color models are fundamental in computer graphics as they define how colors are represented, manipulated, and displayed for digital images.
+In computer graphics applications, understanding and effectively managing color models is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces.
+Color models can be divided into device-oriented models, user-oriented models and device-independent models [32, 33].
+Device-oriented color models are influenced by the signals of the devices and the resulting colors are affected by the technology used for display.
+Widely utilized in various applications that require color consistency with hardware tools, examples include any hardware devices used for human visual perception, such as televisions and video systems.
+Popular color models for devices include RGB, CMY(K) and YCbCr.
+User-oriented color models are considered as a pathway between the observer and the device handling the color information.
+For our study, we mostly investigate cylindrically described models, including HSL and HSV.
+According to [32], these models also inherit other special properties such as device dependence and cylindrical-coordinate representation.
+Device-independent color models are essential for defining color signals without reliance on the specific characteristics of any particular device or application.
+These models play a crucial role in applications that involve color comparisons and the transmission of visual information across networks that connect various hardware platforms.
+We will mostly research color models defined by the International Commission on Illumination (CIE), including CIE XYZ and CIE LAB.
+Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks.
+For example, in [34], a research on traffic light recognition tested a range of color models across different deep learning architectures, with results indicating that the RGB model remains favored despite some of its limitations.
+This suggests that, while RGB is not perfect, it still outperforms other models in certain applications.
+In another study on applying convolutional neural networks (CNNs) for skin detection against adversarial attacks of generative models[35], RGB and YCbCr color models demonstrated superior performance, particularly in handling varying lighting conditions.
+This highlights the adaptability of these color models in scenarios where lighting inconsistencies can degrade model accuracy.
+Moreover, the development of ColorNet[11] illustrates the potential of leveraging multiple color models to capture a broader set of features.
+By combining information from these diverse color models, ColorNet outperforms traditional models that rely on a single color model.
+However, this approach introduces computational overhead due to the inclusion of smaller submodels, which can be a challenge despite the model’s increased efficiency in some cases.
+Despite extensive researches on the use of color models in ANN, their application in SNN remains underexplored and requires further investigation.
+IV Proposed Method
+In this section, we present in detail the design of our coding algorithm.
+IV-A Hypothesis and Preliminary Study
+We hypothesize that the information contained in the bit planes of an image can be effectively used to encode inputs for SNN training.
+Bit planes decompose an image into binary levels, each representing a specific bit of the binary representation of pixel values.
+This decomposition can potentially preserve critical information of the original image while providing a format that might be more suitable for the spike-based processing in SNNs.
 The granularity of information available in bit planes offers a unique approach to encoding visual data, which could enhance the learning capabilities of SNNs by providing a richer, multi-level representation of the input.
-To test our hypotheses, we conducted preliminary experiments using the CIFAR-10 dataset, employing the bit plane encoding method.
-We compared this approach with traditional rate coding and a combination of both methods (illustrated in details in Figure 2 and Section VI-B) with a rate coding step size $T=10$ (illustrated in Figure 1 and Table I).
-Figure 1: Compare training progression on CIFAR-10 [41] between using traditional rate coding (denoted blue), encoding using only bit planes information (denoted orange) and combining bit planes information with rate coding (denoted green).
-Encoding method Validation Accuracy (%)
-Poisson rate coding with time step T=10 70.69
-RGB bit plane encoding 73.49
-Poisson rate coding with time step T=10 and combined RGB bit plane encoding 37.28
-TABLE I: Validation accuracy comparison on CIFAR-10 [41] between traditional rate coding, encoding using only bit planes information and combining bit planes information with rate coding.
-Visible from our results, feeding bit plane information of the encoded RGB color model to the SNNs allowed the model to converge to a certain level of accuracy.
-Still, the accuracy of purely bit planes encoded inputs still fall short compare to rate coding.
-On the other hand, a combination of rate coding and bit planes encoding experience an increase in performance, this outcome demonstrates that bit planes in color models contain valuable information that SNNs can learn from to improve their classification ability and encourage us to perform testing on other datasets with different color models (as mentioned in Section IV-A).
-VI-B Overall encoding scheme with color spaces
-Figure 2: Our proposed encoding method (lower) compare with traditional rate coding method (upper).
-As we can see in Figure 2, for our research the encoding scheme adopted integrates a dual approach featuring both traditional rate coding and bit plane encoding methodologies.
-Initially, the process commences with the conventional rate coding procedure applied to the image tensor.
-Subsequently, depending on the chosen color model (for example RGB), the image tensor may undergo conversion to the specified model or proceed directly to encoding if already in RGB format.
-Following this, the bit plane encoding algorithm (will be discussed in details in Section VI-C and Algorithm 1) is employed to transform the color-modeled image into binary planes.
-The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from rate coding and bit plane encoding.
-These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes, facilitating enhanced performance in our computational models.
-VI-C Bit plane encoding algorithm
-To achieve our research objectives, we have developed Algorithm 1 designed to encode an image tensor batch $X$ into $n_{bit}$ bit planes.
+To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[36], which is a simple dataset containing gray-scale images of hand-written numbers.
+For analysis in term of information contained in the coded spike signals, we perform the calculation of the mean base-2 Shannon entropy[37] $\overline{H}(X_{j})$ as:
+$\overline{H}(X_{j})=-\frac{1}{N}\sum_{j=1}^{N}\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (5)
+and the approximated average Kolmogorov complexity $\overline{K}$ across all dataset’s elements where the Kolmogorov complexity $K(X)$ of a coded image is estimated with DEFLATE compression algorithm[38] $C_{DEFLATE}(X)$ as:
+$\overline{K}(X_{j})$ $=\frac{1}{N}\sum_{j=1}^{N}K(X_{j})$ (6)
+$\approx\frac{1}{N}\sum_{j=1}^{N}C_{DEFLATE}(X_{j})$
+$=\frac{1}{N}\sum_{j=1}^{N}Huffman(LZ77(X_{j}))$
+where $Huffman(X)$ is the Huffman coding[39] and $LZ77(X)$ is the LZ77 compression[40] (or sometimes refered to as sliding window compression).
+We compared three approaches: traditional spike coding, bit plane coding only, and the combination of both spike coding and bit plane coding.
+The detailed settings of this experiment can be found in Section V-A.
+The results are provided in Table I.
+Coding method Validation Accuracy (%) Averaged Kolmogorov complexity approximation (bits) Averaged Shannon entropy (bits)
+Spike coding 98.62 721.5 0.53
+Bit plane coding only 97.87 772.8 0.56
+Combination of spike coding and bit plane coding 98.93 744.4 0.55
+TABLE I: Average validation accuracy and analysis metrics comparison on MNIST[36]
+The results demonstrate that the approximated Kolmogorov complexity and Shannon entropy of bit-plane-coded data are higher, suggesting that the information content is more complex and less predictable.
+This may help explain why the accuracy of models using exclusively bit-plane-encoded inputs is lower compared to the baseline.
+On the other hand, a combination of rate coding and bit-plane coding resulted in an increase in performance.
+This outcome demonstrates that bit planes contain valuable information that SNN can learn from to improve their classification ability.
+This encourages us to carry out a further study and perform testing on other datasets.
+IV-B Proposed Coding Method with Bit Planes
+Figure 2: Our proposed coding method (lower) compare with traditional spike coding method (upper).
+As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding methodologies.
+Initially, the process commences with the conventional spike coding procedure applied to the image tensor.
+The bit plane coding algorithm (Algorithm 1) is employed to transform an input image into binary planes.
+The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from spike coding and bit plane coding.
+These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes.
+Algorithm 1 is designed to encode an image tensor batch $X$ into $n_{bit}$ bit planes.
 To determine $n_{bit}$, it is essential to know the maximum possible value $X_{max}$ of the tensor batch.
-Since different color models have varying $X_{max}$, this value must be predefined according to the selected color model (for example, if RGB is chosen, then $X_{max}=255$).
+Since different color models have different $X_{max}$, this value must be predefined according to the selected color model.
 Once $X_{max}$ is defined, $n_{bit}$ can be calculate as $\left\lceil\log_{2}(X_{max})\right\rceil$.
@@
 This process is repeated until we obtain $n_{bit}$ bit planes.
-The resulting bit planes are then concatenated to form a binary tensor batch with the shape $(n_{bit},B,C,H,W)$.
-This format closely resembles that of rate encoding and ensures determinism in the encoding results without having to sample more than $n_{bit}$ times.
-Algorithm 1 Bit plane encoding algorithm
-$X_{shape}\in(B,C,H,W)$triangleright Constraining input format of multiple images. If input is a single image we should have $B=1$.
-$X_{max}$ is defined
-function BiplaneEncode($X,X_{max}$)
-     $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$.
-     $L\leftarrow[\text{ }]$triangleright Empty list initialization.
+The resulting bit planes are then concatenated to form $n_{bit}$ binary tensor batch of $X_{shape}$.
+This format closely resembles that of rate coding and ensures determinism in the coding results without having to sample more than $n_{bit}$ times.
+Symbol Meaning
+$X$ Input image tensors
+$X_{max}$ Integer value indicate the maximum value of input image tensor
+$n_{bit}$ Minimum number of bit planes required to encode input $X$ into bit planes
+$log_{2}()$ Binary logarithm function
+$X_{shape}$ Dimensions of X, usually include number of images being processed $B$, number of channels $C$ and resolution $H,W$
+$concat()$ Tensor concatenation function, concatenating a list of tensors resulted to a tensor contain spikes of shape ($n_{bit}$, $X_{shape}$)
+$L$ A list-like [41] data structure, used to temporarily store coding result
+$\mathbb{Z}$ The set of all integers
+TABLE II: Table of symbols used in Algorithm 1.
+Algorithm 1 Bit plane coding algorithm
+$X_{max}$ is defined, $X_{max}>0$ and $X\in\mathbb{Z}$
+function BitplaneEncode($X,X_{max}$)
+     $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$
+     $L\leftarrow[\text{ }]$triangleright Empty list initialization
+     for each $X_{shape}.C$ do
          for $i\leftarrow 0,i<n_{bit},i++$ do
-         $L.insert(X\mod{2})$triangleright Element-wise modulo of tensor, then store the bit plane.
-         $X\leftarrow\left\lfloor X/2\right\rfloor$triangleright Element-wise division of tensor.
+              $L.insert(X\mod{2})$triangleright Element-wise modulo of tensor, then store the bit plane to $L$
+              $X\leftarrow\left\lfloor X/2\right\rfloor$triangleright Element-wise division of input image tensor and floor the result
          end for
-     return $concat(L)$triangleright Return a concatenated tensor batches of new dimension $(n_{bit},B,C,H,W)$
+     end for
+     return $concat(L)$triangleright Concatenate all stored sub tensors in $L$ to a single tensor
 end function
-VI-D Experiment setup
-To investigate performance gains during training process, the experimented model should be a medium-sized SNN with potential for improvement in computer vision tasks.
-After careful consideration, we had decided to select SEW-ResNet18 with ADD element-wise function [25] which is a novel medium size SNN tested on various computer vision tasks and suitable to our specifications.
-The model and experiment process was implemented using PyTorch [42] and SpikingJelly[43], incorporating color conversion algorithms based on implementations from Kornia [44].
-During the rate coding steps, all training processes utilized step size $T=10$.
-For optimization, we employed the Adam optimizer [45] with a learning rate of $lr=1e-3$ and $\beta=(0.9,0.999)$ for all tested datasets.
+V Experiments
+V-A Experimental settings
+To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [25].
+The model and experiment process were implemented using PyTorch [42] and SpikingJelly[43].
+Color conversion algorithms are based on implementations from Kornia [44].
+For the baseline method in the experiments, the spike rate coding is employed with a step size of 10.
+To facilitate optimization, we utilize surrogate gradient methods applied to all IF neurons using the arctan surrogate function $S(U)$ [25, 13] with the defined derivative $\partial S$ of:
+$\frac{\partial S}{\partial U}=\frac{1}{\pi}\cdot\frac{1}{1+(\pi U)^{2}}$ (7)
+The optimization process is conducted with the Adam algorithm [45], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets.
 Cross-Entropy Loss with Softmax function and mean reduction $L(x,y)$[42] across $N$ mini-batch dimension was also widely employed as the loss function:
@@
 $l_{n}=-\sum_{c=1}^{C}w_{c}\log\frac{\exp{x_{n,c}}}{\sum_{i=1}^{C}\exp{x_{n,i}}}y_{n,c}$
-with $l_{n}$ is loss of each mini-batch, $x$ is the input, $y$ is the target, $w$ is the weight, total amount of class $C$ where mini-class $c,i\in C$.
-VII Results
-In this section we will discuss results of our main experiments along with some interesting findings beside original hypotheses.
-VII-A The effect of bit plane encoding on colored images
-For our experiment, we conducted tests on various image classification datasets, including CIFAR10[41], CIFAR100[41], EuroSAT[46], Caltech101[47], Caltech256[48], Imagenette[49] (which is a subset of ImageNet[50]), GTSRB[51], and Food101[52].
-We allocated 80% of each dataset for training and the remaining 20% for validation.
-All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU, where CIFAR10[41], CIFAR100[41] and EuroSAT[46] are trained on original input size while input of all other dataset are resize to 224x224 pixels.
-As depicted in Table II, the average accuracy and loss curves across all tested color models demonstrated improved performance.
-Despite a deterministic increase in inference time and input size, the results indicate that the information extracted from color model bit planes enhanced the training process.
-This enhancement serves as an information supplement, leading to higher performance overall without increasing the model size.
-Accuracy of encoding methods (%)
-Dataset Poisson rate coding With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
-CIFAR10[41] 70.69 73.49 73.06 73.64 73.39 73.06 73.63 74.06
-CIFAR100[41] 38.57 42.15 41.31 41.62 41.94 41.91 41.73 42.03
-Caltech101[47] 61.67 64.55 65.65 65.53 66.46 68.41 65.30 66.22
-Caltech256[48] 30.04 41.51 40.65 40.84 40.55 39.94 40.50 40.79
-EuroSAT[46] 84.39 88.48 86.61 88.11 88.74 88.81 88.87 87.48
-Imagenette[49] 72.99 78.68 77.35 78.78 78.55 78.50 77.35 78.70
-Food101[52] 19.62 36.95 37.71 36.98 38.03 36.41 37.10 35.87
-GTSRB[51] 93.36 95.35 51.36 70.68 79.18 80.86 51.78 51.78
-TABLE II: Comparison of best validation accuracy in percentage between traditional rate coding method and our encoding methods, best performing model is emphasized.
-VII-B The effect of bit plane encoding on low resolution grayscale images
-Accuracy of encoding methods (%)
-Dataset Poisson rate coding with time step T=10 Poisson rate coding with time step T=10 and combined RGB bit plane encoding RGB bit plane encoding with no rate coding
-MNIST[53] 98.43 98.51 97.63
-Fashion-MNIST[54] 87.99 88.78 85.76
-KMNIST[55] 93.85 94.60 91.39
-TABLE III: Comparison of best validation accuracy in percentage between encoding methods, best performing model is emphasized.
-Our encoding scheme is designed to leverage color model information to enhance the performance of SNN models, which is why testing on grayscale images was not initially prioritized.
-However, given that bit-plane encoding can be feasibly applied to single-channel grayscale images, we aim to investigate whether this approach yields any significant or insightful results.
-Although grayscale images consist of a pixel intensity range between [0, 255], we can adapt our encoding process by applying RGB-based encoding with a reduced number of channels.
-This method would produce eight bit-plane channels that represent the image information in binary form.
-As demonstrated in Table III (for further details see Figure 5 in Appendix References), the combination of Poisson rate coding with bit-plane encoding outperformed other coding methods in terms of validation accuracy.
-Notably, using only bit-plane information resulted in only a slight reduction in accuracy.
-As discussed in the previous section, we hypothesize that although bit-plane encoding contains less information compared to traditional rate coding, the model can still converge to a certain extent in computer vision tasks.
-We further suspect that this performance deficit may be less pronounced in less complex tasks (in this situation, is a low resolution grayscale image recognition task).
-VIII Conclusion
-In conclusion, our research introduces a novel encoding method for SNN that leverages bit planes derived from various color models of input image data.
+where $l_{n}$ is loss of each mini-batch, $x$ is the input, $y$ is the target, $w$ is the weight, $C$ is the total amount of class where $i\in C$ and $c$ is the mini-classes.
+In the first part of our experiments, we only use grayscale image datasets to clearly see the effects of bit plane coding.
+In the second part, popular datasets of color images are used to see the practical effects of both bit plane coding and color models.
+We allocated 80% of each dataset for training and the remaining 20% with deterministic random seeding for validation.
+All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU.
+V-B Experimental Results
+In this part, we will discuss results of our main experiments along with some interesting findings beside original hypotheses.
+V-B1 The effect of bit plane coding on grayscale images
+Dataset Accuracy of coding methods (%)
+Baseline Bit plane coding only Proposed method
+MNIST[36] 98.62 97.87 98.93
+KMNIST[46] 94.13 92.33 95.58
+Fashion-MNIST[47] 88.83 86.66 90.19
+TABLE III: Comparison of validation accuracy among three coding methods. Best value is highlighted in bold.
+Dataset Analysis results of coding methods (bits)
+Baseline Bit plane coding only Proposed method
+AKC SE AKC SE AKC SE
+MNIST[36] 721.5 0.53 772.8 0.56 744.4 0.55
+KMNIST[46] 1093 0.70 1221 0.70 1150 0.70
+Fashion-MNIST[47] 1384 0.86 1377 0.82 1381 0.84
+TABLE IV: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE).
+Highest values of AKC are denoted in red and lowest are denoted in blue.
+First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information.
+For that, we employed three popular datasets of grayscale images, including MNIST[36], KMNIST[46], Fashion-MNIST[47].
+The results of MNIST dataset were already shown in Table I.
+However, they are still included here to see the behaviors of the coding methods over different datasets.
+As grayscale images consist of only intensity values ranging between [0, 255], we apply our coding process with one channel.
+This process would produce eight bit-plane channels that represent the image information in binary form.
+In this part, we investigate three options, namely 1) the baseline model (spike coding only), 2) using bit plane coding only, 3) the proposed model (spike coding combined with bit plane coding).
+As demonstrated in Table III, the combination of spike coding with bit-plane coding outperformed other coding methods in terms of validation accuracy.
+As mentioned, using only bit-plane coding resulted in a reduction in accuracy; however, the combination of spike coding with bit plane coding (i.e.
+the proposed method) provides gain in all datasets.
+For any method, the accuracy value is highest with the MNIST dataset and lowest with the Fashion-MNIST dataset.
+This could be because among these three datasets, MNIST is the simplest one while Fashion-MNIST is the most complex one.
+Interestingly, the gain of the proposed method compared to the baseline is increased from 0.31% (with MNIST) to roughly 1.4% (with KMNIST and Fashion-MNIST).
+That means the impact of bit plane coding is more prominent with more complex datasets.
+Upon further analysis of spike distribution complexities for various coding methods, as presented in Table IV, several insights emerge.
+First, we observe that the baseline method has the lowest complexity in MNIST and KMNIST datasets, while it has the highest complexity in the Fashion-MNIST dataset.
+We suspect that this difference stems from the nature of these datasets.
+Both MNIST and KMNIST primarily use binary pixel values, with zeros representing the background and ones indicating the presence of a character.
+This binary representation results in a simpler spike distribution.
+In contrast, Fashion-MNIST employs a wider range of pixel values to capture the textures of various clothing items, which introduces greater complexity into the spike distribution when using baseline coding methods.
+This observation suggests that more complex visual datasets may lead to even more pronounced differences in spike distribution when using these coding methods.
+Additionally, by introducing a hybrid coding approach that combines baseline and bit plane coding methods, we hypothesize the potential existence of an optimal complexity threshold at which the SNN might achieve peak performance.
+This idea of an optimizable global maximum suggests that complexity thresholds could play a key role in determining SNN effectiveness.
+Based on our data, we propose that measuring spike distribution complexity could be crucial in tuning SNNs to perform optimally across different datasets and tasks.
+V-B2 The effect of bit plane coding on color images
+Dataset Accuracy of coding methods (%)
+Baseline Bit plane coding only Proposed method
+CIFAR10[26] 70.69 37.28 73.49
+CIFAR100[26] 38.57 11.28 42.15
+Caltech101[48] 61.67 43.28 64.55
+Caltech256[49] 30.04 21.75 41.51
+EuroSAT[50] 84.39 79.87 88.48
+Imagenette[51] 72.99 48.20 78.68
+Food101[52] 19.62 6.29 36.95
+GTSRB[53] 93.36 95.85 95.35
+TABLE V: Average validation accuracy on various computer vision dataset of our proposed method.
+Dataset Analysis results of coding methods (bits)
+Baseline Bit plane coding only Proposed method
+AKC SE AKC SE AKC SE
+CIFAR10[26] 7530 0.995 6764 0.996 7190 0.995
+CIFAR100[26] 7374 0.995 6710 0.996 7079 0.995
+Caltech101[48] 310k 0.993 260k 0.992 279k 0.992
+Caltech256[49] 307k 0.991 257k 0.990 276k 0.990
+EuroSAT[50] 30.3k 0.953 22.9k 0.944 27.0k 0.949
+Imagenette[51] 334k 0.987 282k 0.994 302k 0.992
+Food101[52] 339k 0.989 288k 0.995 308k 0.993
+GTSRB[53] 308k 0.902 2914 0.223 5791 0.019
+TABLE VI: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE), values in thousands are represented with k (e.g., 22k = 22,000).
+Highest values of AKC are denoted in red and lowest are denoted in blue.
+Dataset Analysis results of coding methods (bits)
+With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
+AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE
+CIFAR10[26] 7190 0.995 7086 0.992 6671 0.972 6634 0.927 6752 0.957 7083 0.986 6857 0.988
+CIFAR100[26] 7079 0.995 6975 0.991 6606 0.974 6551 0.929 6660 0.955 6984 0.987 6769 0.985
+Caltech101[48] 279k 0.992 269k 0.978 231k 0.957 241k 0.905 242k 0.929 268k 0.989 239k 0.971
+Caltech256[49] 276k 0.990 265k 0.976 233k 0.957 243k 0.907 244k 0.933 244k 0.989 242k 0.972
+EuroSAT[50] 27.0k 0.949 26.5k 0.968 25.2k 0.944 25.3k 0.993 25.8k 0.950 26.2k 0.932 25.9k 0.966
+Imagenette[51] 302k 0.992 292k 0.989 267k 0.957 271k 0.901 276k 0.941 288k 0.979 277k 0.976
+Food101[52] 308k 0.993 297k 0.990 273k 0.986 277k 0.908 280k 0.927 294k 0.975 281k 0.969
+GTSRB[53] 5791 0.019 5706 0.006 17.6k 0.489 86.1k 0.462 92.9k 0.499 5704 0.005 5703 0.006
+TABLE VII: Comparison of analysis results with our proposed coding method across all tested color models, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE) measured in bits, values in thousands are represented with k (e.g., 22k = 22,000).
+Highest values of AKC are denoted in red and lowest are denoted in blue.
+Dataset Accuracy of coding methods (%)
+With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
+CIFAR10[26] 73.49 73.73 73.64 73.39 73.06 73.63 74.06
+CIFAR100[26] 42.15 41.31 41.62 41.94 41.91 41.73 42.03
+Caltech101[48] 64.55 65.65 65.53 66.46 68.41 65.30 66.22
+Caltech256[49] 41.51 40.65 40.84 40.55 39.94 40.50 40.79
+EuroSAT[50] 88.48 86.61 88.11 88.74 88.81 88.87 87.48
+Imagenette[51] 78.68 77.35 78.78 78.55 78.50 77.35 78.70
+Food101[52] 36.95 37.71 36.98 38.03 36.41 37.10 35.87
+GTSRB[53] 95.35 51.36 70.68 79.18 80.86 51.78 51.78
+Average 65.15 59.21 62.02 63.36 63.49 59.53 59.61
+TABLE VIII: Comparison of validation accuracy among coding methods.
+For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[26], CIFAR100[26], EuroSAT[50], Caltech101[48], Caltech256[49], Imagenette[51] (which is a subset of ImageNet[54]), GTSRB[53], and Food101[52].
+The default color model of these datasets is RGB.
+Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels.
+So, among these datasets, CIFAR10[26], CIFAR100[26] and EuroSAT[50] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
+Across all tested datasets, our proposed methods consistently outperform the baseline method, with the notable exception of the GTSRB[53] dataset, where bit-plane coding alone appears to yield the best performance without the need for combination with traditional spike coding.
+This observation suggests that while our methods generally enhance performance by improving the representation of spike distributions, certain datasets may favor different coding approaches depending on their visual complexity and specific characteristics.
+In the GTSRB[53] case, it seems that bit-plane coding alone is sufficient to capture the essential features for effective processing.
+The analysis results presented in Table VI highlight the relationship between complexity and performance in different coding methods.
+Bit-plane coding consistently exhibits lower complexities across datasets, while traditional spike coding tends to produce higher complexity values.
+This outcome aligns with our hypothesis that complexity increases in visual datasets with greater detail and variation.
+In further investigating the GTSRB[53] dataset, we observed a significant reduction in spike distribution complexity (over 99% when comparing baseline with proposed scheme and pure bit planes coding in Table VI) alongside improved model performance.
+Usually a decrease in complexity can sometimes indicate a loss of critical information, it may also suggest that the spike distribution has become more streamlined, capturing essential features more effectively.
+However, a complexity reduction at such large scale could also introduce elements of instability (shown in Table VII), as it can be challenging to discern whether the decreased complexity represents a beneficial simplification or a detrimental loss of information.
+Consequently, while our findings indicate promising results, we suggest further exploration of complexity thresholds to ensure consistent performance across varying datasets.
+V-B3 Impacts of color models
+In this part, also we investigate two options, namely 1) the baseline method and 2) proposed method.
+The accuracy results of different cases are shown in Table VIII.
+We can see that, when using RGB color model, the accuracy of the proposed method is always higher than that of the baseline.
+The gain ranges from about 2% (with GTSRB[53] dataset) up to 11% (with Caltech256[49] dataset), depending on the employed dataset.
+This suggests that the benefit of the proposed method when applying to color images is more significant than when applying to grayscale images.
+Interestingly, the results show that RGB is not always the best color model.
+Among the eight datasets, RGB provides the best performance in three cases (namely CIFAR100[26], Caltech256[49], GTSRB[53]).
+CMY does not achieve any best results, whereas each of the other five color models achieves the best performance once.
+Especially, with Caltech101[48] dataset, the accuracy of HSV color model is 68.41%, which is much higher than that of RGB color model (64.55%).
+When considering the average of accuracy values across all datasets, the RGB color model is significantly better than other color models (mentioned the last row of Table VIII) and even though there are performance degradation on GTSRB[53], averaged result across all dataset proven to at least out performed baseline method.
+From the above discussion, the key findings of our research can be summarized as follows.
+  • The proposed method provides benefits for both grayscale images and color images.
+  • The gain of the proposed method depends on the employed dataset. For practical color datasets, the gain is from 2% up to 11%.
+  • In terms of average accuracy value over all datasets, RGB is the best color model and CMY is the worst one.
+  • However, considering individual datasets, RGB is not always the best color model. It provides the best performance in only three datasets.
+  • Especially, with the Caltech101 dataset, the performance of HSV color model (68.41%) is much better than that of the RGB color model (64.55%).
+VI Discussion
+VI-A Complexity or entropy?
+In this study, we employ two primary metrics, an approximated Kolmogorov complexity and Shannon entropy, to analyze the effects of bit-plane coding on the distribution of spike trains.
+These metrics were chosen because they are widely used in signal processing and information theory, making them well-suited for analyzing the coded spike and bit-plane signals.
+By treating both coded spikes and bit planes as types of signals, we believe that Kolmogorov complexity and Shannon entropy can provide valuable insights into the complexity of spike distributions.
+While we conducted limited analysis with Shannon entropy, our results indicate a positive correlation between Shannon entropy and Kolmogorov complexity.
+This suggests that Shannon entropy could serve as an alternative to Kolmogorov complexity for those interested in measurement and analysis.
+However, we find Kolmogorov complexity to be more effective, as it provides clearer distinctions between measurements when compared to Shannon entropy.
+Despite achieving substantial insights through these metrics, we recognize that measuring complexity and entropy alone may not fully capture the intricacies of image data.
+Images inherently contain structural complexity and spatial information, which can be difficult to quantify and analyze.
+Focusing solely on spike signals as sequential data may overlook these aspects.
+A helpful analogy is the historical debate in physics regarding the nature of light, with some scientists viewing light as a wave and others as particles.
+It was only with the discovery of wave-particle duality through experiments like the double-slit experiment that a more comprehensive understanding emerged.
+Similarly, viewing spike signals solely as sequential signals could reveal certain patterns but may not capture the full complexity underlying the optimization process in image processing tasks.
+Currently, tools designed specifically for spike signal analysis that incorporate spatial information for image-based tasks are underexplored.
+This gap highlights the need for further research into analysis methods that can integrate both signal and spatial information for more complete insights.
+VII Conclusion
+In conclusion, our research introduces a novel coding method for SNN that leverages bit planes derived from various color models of input image data.
 This approach prove to enhances the computational accuracy of SNN without increasing the model size.
-Through extensive experimental validation, we have demonstrated the effectiveness of our encoding strategy across multiple computer vision tasks, achieving notable performance gains compared to conventional methods.
-This study is pioneering in its application of color spaces within the context of spiking neural networks, and we believe that our findings open new avenues for the development of more efficient and effective SNN models.
+Through extensive experimental validation, we have demonstrated the effectiveness of our coding strategy across multiple computer vision tasks, achieving notable performance gains compared to conventional methods.
+This study is pioneering in its application of color models within the context of spiking neural networks, and we believe that our findings open new avenues for the development of more efficient and effective SNN models.
 Our work sets a foundation for future research and applications, potentially advancing the field of computational neuroscience and artificial intelligence.
+Appendix A CMY or CMYK?
+In the initial stages of our experimental setup and algorithm design, we encountered challenges in selecting the appropriate device-dependent color model for our experiments.
+Specifically, we debated whether to use the CMY or CMYK color model.
+Each has advantages in different applications, but their compatibility with the neural networks used in our research required careful consideration.
+In image processing and printing, the CMYK model was developed to improve print quality by reducing ink usage in darker regions.
+This is achieved by adding a K component, which reduces the CMY values proportionally, leading to enhanced contrast and more economical ink usage in printed images.
+However, when adapting this approach for digital image processing, particularly for direct usage on digital devices, the inclusion of a K channel becomes less beneficial.
+Unlike printed images, digital displays do not require ink density adjustments, making the K channel unnecessary for accuracy.
+For our experiment, this would introduced dimensional conflicts in tested SNN architectures and coding schemes since they are previously optimized for three-channel data inputs (RGB), including a fourth channel would require additional modifications to model architecture, potentially affecting performance and comparability.
+Furthermore, in our image processing pipeline, we observed that converting images to the CMYK color model using standard image processing libraries (e.g., PIL[55]) did not apply undercolor removal during conversion to CMYK model, resulting in a K channel of zero values.
+This posed a unique problem in our training setup: a zeroed K channel in the input data could nullify both the input signal and its gradients when surrogate gradients are used for backpropagation in SNNs.
+Consequently, this would undermine the learning process by reducing the effectiveness of gradient-based updates, complicating model training and ultimately impacting the accuracy of our results.
+Appendix B Tranning progession details
+This section contain the training progression graph for each variant of experiments in Section V-B3 and Section V-B1.
+It is noteworthy from Figure 4 and 5 that performance fluctuations were observed in both traditional coding and our coding method.
+We hypothesize that these fluctuations might related to inherent error rate during sampling process in Poisson-distributed spike trains[3].
+Another hypothesis is that although bit planes may possess a certain degree of spatial information, they can also encompass noise on low order bit planes that interferes with the training process or have little to none distinguishable features (which might explain how on low noises single-channel image dataset such as MNIST, the performance of only using bit planes coding remain competitive to traditional rate coding with less amount of time step).
+Consequently, the selective identification of bit planes that retain valuable spatial information continues to pose a significant challenge.
+(a) MNIST
+(b) Fashion-MNIST
+(c) KMNIST
+Figure 3: Comparison in accuracy progression and loss progression between coding methods on different low-resolution gray-scale image classification datasets.
+(a) CIFAR10[26]
+(b) CIFAR100[26]
+(c) Caltech101[48]
+(d) Caltech256[49]
+(e) EuroSAT[50]
+(f) Imagenette[51]
+(g) Food101[52]
+(h) GTSRB[53]
+Figure 4: Comparison in loss progression of traditional rate coding with our method on image classification datasets.
+On every sub-figure, loss details of every color models is shown on the left and average result of all models on the right with maximum loss difference denoted on top of figure.
+(a) CIFAR10[26]
+(b) CIFAR100[26]
+(c) Caltech101[48]
+(d) Caltech256[49]
+(e) EuroSAT[50]
+(f) Imagenette[51]
+(g) Food101[52]
+(h) GTSRB[53]
+Figure 5: Comparison in accuracy progression of traditional rate coding with our method on image classification datasets.
+On every sub-figure, accuracy details of every color models is shown on the left and average result of all models on the right with maximum accuracy difference denoted on top of figure.

### v3

@@
 However, their practical adoption is constrained by several challenges, prominently among them being performance optimization.
-In this study, we present a novel approach to enhance the performance of SNN for images through a new coding method that exploits bit plane representation.
-Our proposed technique is designed to improve the accuracy of SNN without increasing model size.
+In this study, we present a novel approach to enhance the SNN for image classification through a new coding method that exploits bit plane representation.
 Also, we investigate the impacts of color models of the proposed coding process.
-Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets.
+Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets of grayscale images and color images.
 To the best of our knowledge, this is the first research that considers bit planes and color models in the context of SNN.
-By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more efficient and effective SNNs models in future researches and applications.
+By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more effective SNNs models in future researches and applications.
+Source code for our experiments is available at https://github.com/luutn2002/bit-plane-snn
 Index Terms:
-Spiking neural network, spike coding, image classification
+Spiking neural network, spike coding, image classification, bit planes, color models
 I Introduction
@@
 This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[4].
-The inherent sparsity and temporal coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [5].
+The inherent sparsity and temporality in coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [5].
 Despite their potential, SNNs pose significant challenges, particularly in training methodologies and network optimization.
 Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[6].
-Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9], have been developed to address these challenges, enabling the effective training of SNNs.
+Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9, 10, 11], have been developed to address these challenges, enabling the effective training of SNNs.
 Prior research in ANN has explored the application of bit planes (e.g.
@@
 This led us to formulate our research questions:
-  • Does bit plane representation enhance the performance of SNN?
-  • What are impacts of different color models on SNN?
-Especially, we observe that bit planes of an image contain pulse-like information, and can be applied to encode color information for SNNs to possibly improve accuracy.// After extensive experimentation, we concluded that bit planes contain valuable information that can be utilized in the rate coding process as a supplementary source to enhance the accuracy of direct surrogate backpropagation process [13] in SNN architectures.
-Furthermore, bit plane information proved to be reusable across multiple color models in most of the cases examined in our study.
-A preliminary version of this paper is published in [14], where the benefit of including bit plane coding is proved by three grayscale datasets.
+  • How can we utilize bit plane representation to enhance the performance of SNN?
+  • How can we utilize different color models to enhance the performance of SNN?
+In our research, we observe that bit planes of an image contain pulse-like information, and can be applied to encode pixel values for SNNs to improve accuracy during surrogate gradient training.
+Through extensive experiments, it is shown that bit planes contain valuable information that can be utilized in the spike coding process as a supplementary data source for SNN.
+Also, bit plane coding proves to be reutilized across multiple color models.
+Furthermore, RGB color model, which is the default color model of many datasets, is not always the best one among the evaluated color models in our study.
+A preliminary version of this paper is published in [15], where the benefit of using bit plane coding is proved by three grayscale datasets.
 In this extended version, the following new points have been added.
   • First, the proposed method was updated to support color images.
-Second, eight more datasets are included.
-Third, impacts of color models are investigated.
+  • Second, eight more datasets were included in the experiments.
+  • Third, impacts of color models were investigated.
+  • Fourth, discussion and analysis of the paper were significantly extended.
+TABLE I: Table of abbreviation (in alphabetical order) used in this paper.
+Abbreviation Full Name
+ADD Addition
+ANN Artificial neural network
+CIE International Commission on Illumination
+IF Integrate-and-fire
+LIF Leaky integrate-and-fire
+ReLU Rectified linear unit
+SEW Spike-element-wise
+SNN Spiking neural network
+SNR Signal-to-noise ratio
+STDP Spike-timing-dependent plasticity
+TABLE II: Table of algorithmic symbols used in this paper.
+Symbol Meaning
+$X$ Input image tensors
+$X_{max}$ Integer value indicate the maximum value of input image tensor
+$n_{bit}$ Minimum number of bit planes required to encode input $X$ into bit planes
+$log_{2}()$ Binary logarithm function
+$X_{shape}$ Dimensions of X, include: $B$: number of images being processed. $C$: number of image channels. $H$: image height. $H$: image width.
+$concat()$ Tensor concatenation function, concatenating a list of tensors resulted to a tensor contain spikes of shape ($n_{bit}$, $X_{shape}$)
+$convertColor()$ Tensor color model converting function, convert $X$ to a specified color model decided by value of $colormodel$
+$colormodel$ A string value specify the color model to convert the tensor to (i.e. ”RGB”, ”CMY”)
+$mod$ Modulo division
+$L$ A list-like [16] data structure, used to temporarily store coding result
+$\mathbb{Z}$ The set of all integers
 II Spiking Neural Network
-In this section we present an overview of SNN knowledge related to our research.
-II-A History of SNNs
-Phenomenological neuron models can be progressively simplified until reaching a stage where even neural spikes are not modeled, leading to first and second generation neurons [15, 16].
-This simplification significantly reduces computational demands but also decreases biological realism.
-The detail level in modeling often correlates with available computing power.
-For example, ANNs based on second generation neurons were historically limited by computing power [17], restricting the size of networks and the complexity of neuronal models.
-This limitation influenced the design of learning algorithms.
-Even with advancements in computing power, the complexity of neuronal models did not proportionately increase because existing learning algorithms were not compatible with more detailed models.
-Consequently, two distinct research fields emerged.
-  • ANNs[18] focused on large networks of simplified neuron-like units, capable of learning through algorithms like backpropagation, and were successfully applied in pattern recognition, classification, and pattern completion tasks.
-  • Computational Neuroscience[19] in contrast used detailed biophysical and phenomenological models in smaller networks to study electrophysiological processes, pattern generation, and dynamic behaviors of small neuron groups.
-Although there have been attempts to create large networks of detailed biophysical neuron models [20], it remains challenging due to computational constraints.
-Recent advances in computing power have blurred the lines between these fields.
-ANNs have become more biologically realistic, while networks of biophysical neurons have grown larger and more detailed.
-However, computational power still limits the application of detailed models in large neural networks for pattern recognition and classification.
-With improved computing power, new learning algorithms are being developed for these models.
-SNNs, which use spiking neurons modeled by phenomenological models like the Spike Response Model[21], represent the next step towards this goal.
-Due to the computational burden, biophysical models are less common in certain SNN applications.
-This paper focuses on research involving SNNs with phenomenological models, excluding those using biophysical models.
-II-B The Integrate-and-Fire model
-One of the earliest neuron models is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [22].
+In this section we present an overview of SNN. The abbreviations and algorithmic symbols used throughout the paper are listed in Table I and Table II.
+II-A Overview of SNNs
+SNNs can be considered a biologically inspired variant of ANNs that offer a different computational paradigm.
+Unlike traditional ANNs, which process information in a continuous manner using floating-point activations and backpropagation-based learning [17], SNNs rely on discrete events called spikes to encode and transmit information [18].
+This event-driven nature of SNNs allows for energy-efficient computations, making them suitable for neuromorphic hardware implementations [19].
+While the majority of mainstream ANNs architecture operate on floating point data representations, SNNs mimic the sparse, time-dependent processing found in biological neurons [20].
+Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [21].
+Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [9] or ANN-to-SNN conversion approaches [22].
+Thus, SNNs present a compelling alternative to ANNs for energy-efficient and biologically plausible computations, while also introducing unique challenges in model design and training.
+One of the earliest neuron models that is still widely applied in modern SNN architectures is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [23].
 In this model, the neuron is characterized by its membrane potential $V$:
@@
 In this model, the firing rate increases linearly and indefinitely with the increase in input current.
-While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons, challenges still remain in the conversion process.
+While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons[24, 25, 26], challenges still remain in the conversion process.
 One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches.
@@
 This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [28].
-Still, within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better LIF neurons[24], with the SEW-ResNet[25] architecture being one such example.
-Figure 1: Illustrated bit planes of a sample image from CIFAR10[26] dataset.
-II-C Deep Residual Learning in Spiking Neural Networks
+Still, recent study have shown that SNN employed with IF neurons are more energy efficient than traditional ANN[29].
+Within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better than leaky integrate-and-fire (LIF) neurons[28], with the SEW-ResNet[11] architecture being one such example.
+Figure 1: Visualization of the extracted bit planes from a sample image in the CIFAR-10 dataset [30].
+Higher-order bit planes (e.g., most significant bits) encode more spatial and structural information, capturing the prominent features of the image.
+In contrast, lower-order bit planes (e.g., least significant bits) primarily contain finer details and noise, contributing less to the overall image structure.
+II-B Deep Residual Learning in SNNs
 The success of ResNet[31] in deep learning has motivated efforts to extend residual learning to SNN.
@@
 While this is straightforward in ResNets[31] with ReLU activation, it is not always feasible in Spiking ResNets[26].
-Specifically, [25] claimed that for certain spiking neuron models, such as leaky integrate-and-fire (LIF) neurons[22] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging.
+Specifically, [11] claimed that for certain spiking neuron models, such as LIF neurons[23] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging.
 This complicates the realization of identity mapping.
@@
 This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models.
-SEW ResNet [25] was also the first work to propose and apply a surrogate gradient approach for optimizing SNNs.
-Using backpropagation methods to train SNNs [29, 13] has become a common practice in neuromorphic computing.
-However, applying backpropagation in SNNs presents certain challenges.
+In recent researches, using backpropagation methods to train SNNs [32, 10, 9, 11] has become a common practice in neuromorphic computing.
 In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
@@
 As shown, directly differentiating $H(\theta,t)$ results in unstable gradients.
-To address this, [25] proposed substituting the derivative of the Heaviside function $\partial H$ with a numerically stable gradient from a different activation function during backpropagation, such as the derivative $\partial\sigma$ of the Sigmoid function $\sigma(\theta,t)$:
+To address this, we can substitute the derivative of the Heaviside function $\partial H$ with a numerically stable gradient from a different activation function during backpropagation, such as the derivative $\partial\sigma$ of the Sigmoid function $\sigma(\theta,t)$:
 $\sigma(\theta,t)=\frac{1}{1+e^{\theta-t}}$ (4)
 $\frac{\partial H}{\partial\theta}=\frac{\partial\sigma}{\partial\theta}=\frac{e^{\theta-t}}{(e^{\theta-t}+1)^{2}}$
-This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively.
-With these advantages, we decided to choose SEW Resnet[25] architecture for our study.
+This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively [11].
+Similar to other new types of neural networks [33, 34, 35], the choice of input data representation may have strong impacts on output performance.
+Thus, in this paper we investigate a new way that employs bit plane coding for data representation in SNNs.
 III Bit Planes and Color Models in Deep Learning
@@
 III-A Bit planes in deep learning
+In the context of digital discrete signals such as images, a bit plane represents the collection of bits that occupy the same position within the binary representation of the signal’s numerical values (as illustrated in Figure 1).
+For instance, in a 8-bit data representation, there are 8 distinct bit planes.
+The $1^{st}$ bit plane consists of the most significant bits, while the $8^{th}$ bit plane comprises the least significant bits.
+The first bit plane provides the coarsest yet most fundamental approximation of the signal’s values.
+As the bit plane index increases, its contribution to the overall representation diminishes.
+Consequently, incorporating additional bit planes progressively refines the signal’s approximation.
 Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance.
@@
 We will mostly research color models defined by the International Commission on Illumination (CIE), including CIE XYZ and CIE LAB.
-Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks.
-For example, in [34], a research on traffic light recognition tested a range of color models across different deep learning architectures, with results indicating that the RGB model remains favored despite some of its limitations.
-This suggests that, while RGB is not perfect, it still outperforms other models in certain applications.
-In another study on applying convolutional neural networks (CNNs) for skin detection against adversarial attacks of generative models[35], RGB and YCbCr color models demonstrated superior performance, particularly in handling varying lighting conditions.
-This highlights the adaptability of these color models in scenarios where lighting inconsistencies can degrade model accuracy.
-Moreover, the development of ColorNet[11] illustrates the potential of leveraging multiple color models to capture a broader set of features.
-By combining information from these diverse color models, ColorNet outperforms traditional models that rely on a single color model.
-However, this approach introduces computational overhead due to the inclusion of smaller submodels, which can be a challenge despite the model’s increased efficiency in some cases.
-Despite extensive researches on the use of color models in ANN, their application in SNN remains underexplored and requires further investigation.
+Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks, such as image recognition[40, 13], image segmentation[14] and adversarial attacks prevention[41].
+Despite extensive researches on the use of color models in ANN, their application in SNN remains unexplored and requires further investigation.
+Figure 2: Our proposed coding method (lower) compare with traditional spike coding method (upper).
 IV Proposed Method
 In this section, we present in detail the design of our coding algorithm.
-IV-A Hypothesis and Preliminary Study
 We hypothesize that the information contained in the bit planes of an image can be effectively used to encode inputs for SNN training.
@@
 The granularity of information available in bit planes offers a unique approach to encoding visual data, which could enhance the learning capabilities of SNNs by providing a richer, multi-level representation of the input.
-To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[36], which is a simple dataset containing gray-scale images of hand-written numbers.
-For analysis in term of information contained in the coded spike signals, we perform the calculation of the mean base-2 Shannon entropy[37] $\overline{H}(X_{j})$ as:
-$\overline{H}(X_{j})=-\frac{1}{N}\sum_{j=1}^{N}\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (5)
-and the approximated average Kolmogorov complexity $\overline{K}$ across all dataset’s elements where the Kolmogorov complexity $K(X)$ of a coded image is estimated with DEFLATE compression algorithm[38] $C_{DEFLATE}(X)$ as:
-$\overline{K}(X_{j})$ $=\frac{1}{N}\sum_{j=1}^{N}K(X_{j})$ (6)
-$\approx\frac{1}{N}\sum_{j=1}^{N}C_{DEFLATE}(X_{j})$
-$=\frac{1}{N}\sum_{j=1}^{N}Huffman(LZ77(X_{j}))$
-where $Huffman(X)$ is the Huffman coding[39] and $LZ77(X)$ is the LZ77 compression[40] (or sometimes refered to as sliding window compression).
-We compared three approaches: traditional spike coding, bit plane coding only, and the combination of both spike coding and bit plane coding.
-The detailed settings of this experiment can be found in Section V-A.
-The results are provided in Table I.
-Coding method Validation Accuracy (%) Averaged Kolmogorov complexity approximation (bits) Averaged Shannon entropy (bits)
-Spike coding 98.62 721.5 0.53
-Bit plane coding only 97.87 772.8 0.56
-Combination of spike coding and bit plane coding 98.93 744.4 0.55
-TABLE I: Average validation accuracy and analysis metrics comparison on MNIST[36]
-The results demonstrate that the approximated Kolmogorov complexity and Shannon entropy of bit-plane-coded data are higher, suggesting that the information content is more complex and less predictable.
-This may help explain why the accuracy of models using exclusively bit-plane-encoded inputs is lower compared to the baseline.
-On the other hand, a combination of rate coding and bit-plane coding resulted in an increase in performance.
+To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[42], which is a simple dataset containing grayscale images of hand-written numbers.
+We compared three approaches: traditional rate coding, bit plane coding only, and the combination of both rate coding and bit plane coding.
+The detailed settings of this experiment can be found in Section VI-A.
+The results are provided in Table III.
+TABLE III: Average validation accuracy comparison on MNIST[42]
+Coding method Validation Accuracy (%)
+Rate coding 98.62
+Bit plane coding only 97.87
+Combination of rate coding and bit plane coding 98.93
+We can see that the accuracy when using exclusively bit-plane-encoded inputs is lower compared to the baseline.
+However, a combination of rate coding and bit-plane coding resulted in an increase in performance.
 This outcome demonstrates that bit planes contain valuable information that SNN can learn from to improve their classification ability.
 This encourages us to carry out a further study and perform testing on other datasets.
-IV-B Proposed Coding Method with Bit Planes
-Figure 2: Our proposed coding method (lower) compare with traditional spike coding method (upper).
-As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding methodologies.
+As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding.
 Initially, the process commences with the conventional spike coding procedure applied to the image tensor.
-The bit plane coding algorithm (Algorithm 1) is employed to transform an input image into binary planes.
+The bit plane coding algorithm is then employed to transform an input image into binary planes.
 The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from spike coding and bit plane coding.
 These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes.
-Algorithm 1 is designed to encode an image tensor batch $X$ into $n_{bit}$ bit planes.
-To determine $n_{bit}$, it is essential to know the maximum possible value $X_{max}$ of the tensor batch.
-Since different color models have different $X_{max}$, this value must be predefined according to the selected color model.
-Once $X_{max}$ is defined, $n_{bit}$ can be calculate as $\left\lceil\log_{2}(X_{max})\right\rceil$.
-To calculate the bit planes of an unnormalized integer input tensor batch $X$, we perform element-wise division of $X$ by 2.
-The remainder from this division constitutes a bit plane of $X$, while the quotient is saved for the extraction of subsequent bit planes.
-This process is repeated until we obtain $n_{bit}$ bit planes.
-The resulting bit planes are then concatenated to form $n_{bit}$ binary tensor batch of $X_{shape}$.
-This format closely resembles that of rate coding and ensures determinism in the coding results without having to sample more than $n_{bit}$ times.
-Symbol Meaning
-$X$ Input image tensors
-$X_{max}$ Integer value indicate the maximum value of input image tensor
-$n_{bit}$ Minimum number of bit planes required to encode input $X$ into bit planes
-$log_{2}()$ Binary logarithm function
-$X_{shape}$ Dimensions of X, usually include number of images being processed $B$, number of channels $C$ and resolution $H,W$
-$concat()$ Tensor concatenation function, concatenating a list of tensors resulted to a tensor contain spikes of shape ($n_{bit}$, $X_{shape}$)
-$L$ A list-like [41] data structure, used to temporarily store coding result
-$\mathbb{Z}$ The set of all integers
-TABLE II: Table of symbols used in Algorithm 1.
+The process of converting a multi-channel image into its bit planes involves isolating each bit of the binary representation of the pixel values for each color channel.
+For an $n_{bit}$ image, let $I_{C}(x,y)$ denote the intensity of a pixel at position $(x,y)$ for a given color channel $C$, pixel intensity can be expressed in terms of its binary representation as:
+$I_{C}(x,y)=\sum_{k=0}^{n_{bit}}B_{k,C}(x,y)\cdot 2^{k}$ (5)
+where $B_{k,C}(x,y)\in\{0,1\}$ is the $k$-th bit of the binary representation of $I(x,y)$, and $k$ represents the bit-plane index, with $k=0$ corresponding to the least significant bit and $k=n_{bit}$ corresponding to the most significant bit.
+To isolate a specific bit-plane, the $k$-th bit of the pixel can be extracted using the following equation:
+$B_{k,C}(x,y)=\left\lfloor\frac{I_{C}(x,y)}{2^{k}}\right\rfloor\mod 2$ (6)
 Algorithm 1 Bit plane coding algorithm
 $X_{max}$ is defined, $X_{max}>0$ and $X\in\mathbb{Z}$
-function BitplaneEncode($X,X_{max}$)
+function BitplaneEncode($X,X_{max}$, $colormodel$)
+     $X=convertColor(X,colormodel)$triangleright Convert $X$ to specified color model
      $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$
@@
 end function
-V Experiments
-V-A Experimental settings
+To encode an image tensor batch $X$ into $n_{bit}$ bit planes in our training process based on established knowledge, we implement Algorithm 1.
+It is essential to know the maximum possible value $X_{max}$ of the tensor batch in order to determine $n_{bit}$.
+Since different color models have different $X_{max}$, this value must be predefined according to the selected color model.
+Once $X_{max}$ is defined, $n_{bit}$ can be calculate as:
+$n_{bit}=\lceil\log_{2}(X_{max})\rceil$ (7)
+Bit planes calculation of an unnormalized integer input tensor batch $X$ can be performed by element-wise division of $X$ by 2.
+The remainder from this division constitutes a bit plane of $X$, while the quotient is saved for the extraction of subsequent bit planes.
+This process is repeated until we obtain $n_{bit}$ bit planes.
+The resulting bit planes are then concatenated to form $n_{bit}$ binary tensor batch of $X_{shape}$.
+This format closely resembles that of rate coding and ensures determinism in the coding results without having to sample more than $n_{bit}$ times.
+V Theoretical basis
+In this part, we aim to establish a theoretical foundation for our proposed method of spike coding within surrogate gradient approaches.
+As discussed in Section II-B, surrogate gradient methods attempt to adapt deep learning’s gradient optimization techniques to SNNs, maintaining a close connection to traditional deep learning and convex optimization.
+A classical stochastic optimization problem associated with learning a target function $h$ can be formulated as[43]:
+$\min_{\mathbf{w}}F_{h}(\mathbf{w})=\mathbb{E}_{\mathbf{x}}\left[\ell(p_{\mathbf{w}}(\mathbf{x}),h(\mathbf{x}))\right],$ (8)
+where $\ell$ is a loss function, $\mathbf{x}$ represents stochastic inputs (assumed to be vectors in Euclidean space), and $p_{\mathbf{w}}$ is a differentiable predictor parameterized by $\mathbf{w}$.
+It is assumed that $E_{x}[\left\|\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}\right\|^{2}]\leq G(\mathbf{w})^{2}$ for some scalar $G(\mathbf{w})$, and that $F$ is differentiable[43, 44].
+By reformulating our surrogate gradient optimization process as a traditional stochastic optimization problem, our objective can be defined as:
+$\min_{\mathbf{w}}F_{h}(\mathbf{w})$ $=\mathbb{E}_{\mathbf{x}}\left[\ell(P_{\mathbf{w}}(\mathbf{x}),h(\mathbf{x}))\right],$ (9)
+$P_{\mathbf{w}}(\mathbf{x})$ $=\mathbb{E}_{T}\left[p_{\mathbf{w}}(e(\mathbf{x}))\right],$
+where the surrogate gradient model $P_{\mathbf{w}}(\mathbf{x})$ encodes the input $\mathbf{x}$ into a spike signal using an encoding function $e(\mathbf{x})$ and computes the expected output signal across the time dimension $T$ to generate the final stochastic prediction.
+Previous studies in gradient-based optimization[43, 44] have demonstrated a positive correlation between a model’s performance and the signal-to-noise ratio (SNR) of the stochastic gradient, defined by:
+$SNR(\delta(\mathbf{x}))=\frac{Sig(\delta(\mathbf{x}))}{Noi(\delta(\mathbf{x}))},$ (10)
+$Sig(\delta(\mathbf{x}))=\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}$
+$Noi(\delta(\mathbf{x}))=\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}$
+$\delta(\mathbf{x})=h(\mathbf{x})g_{\mathbf{w}}(\mathbf{x})$
+where $g_{{\mathbf{w}}}({\mathbf{x}})=\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}({\mathbf{x}})$ is the Jacobian of model $p_{{\mathbf{w}}}({\mathbf{x}})$ with respect to model weight $\mathbf{w}$.
+Since this correlation does not depend on input gradient calculation, we can directly apply this in our research.
+To improve the SNR, we can either increase signal term $Sig(\delta(\mathbf{x}))$ or reduce noise term $Noi(\delta(\mathbf{x}))$.
+As shown in [11], the norm $\left\|\delta({\mathbf{x}})\right\|$ is positively correlated with timestep $T$:
+$\left\|\delta({\mathbf{x}})\right\|=\left\|h^{\prime}({\mathbf{x}})\right\|\cdot\left\|\frac{\partial\ell}{\partial O^{l+k-1}}\right\|\cdot\left\|\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)\right\|$ (11)
+$\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)=\sqrt{NT\alpha}$
+$\alpha=\left[\Phi\,\Theta^{\prime}(1-V_{th})^{2k}+(1-\Phi)\,\Theta^{\prime}(0-V_{th})^{2k}\right]$
+$\Phi=\frac{\sum_{j=0}^{N-1}\sum_{t=0}^{T-1}S^{l}_{j}[t]}{N\,T}$
+where the equation denote the magnitude of $\left\|\delta({\mathbf{x}})\right\|$ with target function gradient $h^{\prime}({\mathbf{x}})$, previous blocks’ gradient $\frac{\partial L}{\partial O^{l+k-1}}$ of a $k$ sequential blocks $l$ layers Spiking-ResNet architecture[11] and Heaviside function’s gradient $\Theta^{\prime}$ while transmitting signal $S$ of previous layer output $O$ at firing rate $\Phi$ using $N$ neurons with threshold voltage $V_{th}$.
+Upon performing signal concatenation using bit-plane-based signal $B$ of bit length $N_{bit}$ as in the proposed method, the magnitude of $\left\|\delta({\mathbf{x}})\right\|$ and firing rate $\Phi$ would increase as:
+[…]
+$\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)\Theta^{\prime}\left(B^{l+i}-V_{th}\right)=\sqrt{N(T+N_{bit})\alpha}$
+$\alpha=\left[\Phi\,\Theta^{\prime}(1-V_{th})^{2k}+(1-\Phi)\,\Theta^{\prime}(0-V_{th})^{2k}\right]$
+$\Phi=\frac{\sum_{j=0}^{N-1}\sum_{t=0}^{T-1}S^{l}_{j}[t]+\sum_{j=0}^{N-1}\sum_{t=0}^{N_{bit}-1}B^{l}_{j}[t]}{N\,(T+N_{bit})}$
+and improving the chance of obtaining higher $SNR(\delta(\mathbf{x}))$.
+Figure 3: Comparing model gradient’s SNR of the proposed method (red), baseline (blue) and bit plane coding (green) of MNIST dataset across the epochs.
+Dashed line indicate the mean gradient SNR (higher is better based on our assumption).
+To empirically validate our hypothesis, we measure the gradient SNR of the model across all training epochs on MNIST dataset (since to ensure that gradient does not saturate during training, $p_{\mathbf{w}}({\mathbf{x}})$ should serves as a reasonable approximation of the target function $h({\mathbf{x}})$).
+The resulted SNR measurements across all training epochs are presented in Figure 3.
+The findings demonstrate that our proposed method achieves a higher mean SNR during training process, which aligns with our initial assumption.
+VI Experiments
+VI-A Experimental settings
 To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [11].
@@
 Color conversion algorithms are based on implementations from Kornia [47].
-For the baseline method in the experiments, the spike rate coding is employed with a step size of 10.
-To facilitate optimization, we utilize surrogate gradient methods applied to all IF neurons using the arctan surrogate function $S(U)$ [25, 13] with the defined derivative $\partial S$ of:
-$\frac{\partial S}{\partial U}=\frac{1}{\pi}\cdot\frac{1}{1+(\pi U)^{2}}$ (7)
-The optimization process is conducted with the Adam algorithm [45], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets.
-Cross-Entropy Loss with Softmax function and mean reduction $L(x,y)$[42] across $N$ mini-batch dimension was also widely employed as the loss function:
-$L(x,y)=\frac{\sum_{n=1}^{N}l_{n}}{N},$ (8)
-$l_{n}=-\sum_{c=1}^{C}w_{c}\log\frac{\exp{x_{n,c}}}{\sum_{i=1}^{C}\exp{x_{n,i}}}y_{n,c}$
-where $l_{n}$ is loss of each mini-batch, $x$ is the input, $y$ is the target, $w$ is the weight, $C$ is the total amount of class where $i\in C$ and $c$ is the mini-classes.
+For the baseline method in the experiments, the Poisson-distributed rate coding [48, 49] is employed with a step size $T=10$ where spike $s_{T,i}$ for i-th pixel at step $T$ with intensity $P_{T,i}$ can be interpreted as:
+$s_{T,i}=\begin{cases}1&\text{if }random(0.0,1.0)\geq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (13)
+To enable efficient training, surrogate gradient methods were applied to all IF neurons.
+We utilized the arctangent surrogate gradient function $S(U)$[11, 10].
+Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[45, 50, 51], computed using mean reduction across mini-batches.
+Inspired by traditional deep learning approach in model fine-tunning [50, 51], the optimization process is conducted with the Adam algorithm [52], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets.
 In the first part of our experiments, we only use grayscale image datasets to clearly see the effects of bit plane coding.
@@
 All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU.
-V-B Experimental Results
+VI-B Experimental Results
 In this part, we will discuss results of our main experiments along with some interesting findings beside original hypotheses.
-V-B1 The effect of bit plane coding on grayscale images
+VI-B1 The effect of bit plane coding on grayscale image datasets
+TABLE IV: Comparison of validation accuracy among coding methods. Best performances on each dataset are denoted in bold. Performance gain when compared with the baseline is highlighted in green.
 Dataset Accuracy of coding methods (%)
-Baseline Bit plane coding only Proposed method
-MNIST[36] 98.62 97.87 98.93
-KMNIST[46] 94.13 92.33 95.58
-Fashion-MNIST[47] 88.83 86.66 90.19
-TABLE III: Comparison of validation accuracy among three coding methods. Best value is highlighted in bold.
-Dataset Analysis results of coding methods (bits)
-Baseline Bit plane coding only Proposed method
-AKC SE AKC SE AKC SE
-MNIST[36] 721.5 0.53 772.8 0.56 744.4 0.55
-KMNIST[46] 1093 0.70 1221 0.70 1150 0.70
-Fashion-MNIST[47] 1384 0.86 1377 0.82 1381 0.84
-TABLE IV: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE).
-Highest values of AKC are denoted in red and lowest are denoted in blue.
+Baseline Bit plane coding Proposed method
+MNIST[42] 98.62 97.87 98.93 (+0.31)
+KMNIST[53] 94.13 92.33 95.58 (+1.45)
+Fashion-MNIST[54] 88.83 86.66 90.19 (+1.36)
+Average 93.86 92.29 94.90 (+1.04)
 First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information.
 For that, we employed three popular datasets of grayscale images, including MNIST[42], KMNIST[53], Fashion-MNIST[54].
-The results of MNIST dataset were already shown in Table I.
+The results of MNIST dataset were already shown in Table III.
 However, they are still included here to see the behaviors of the coding methods over different datasets.
 As grayscale images consist of only intensity values ranging between [0, 255], we apply our coding process with one channel.
-This process would produce eight bit-plane channels that represent the image information in binary form.
-In this part, we investigate three options, namely 1) the baseline model (spike coding only), 2) using bit plane coding only, 3) the proposed model (spike coding combined with bit plane coding).
-As demonstrated in Table III, the combination of spike coding with bit-plane coding outperformed other coding methods in terms of validation accuracy.
+This process would produce 8 bit-plane channels that represent the image information in binary form.
+In this part, we investigate three options,namely:
+  • the baseline model (spike coding only),
+  • using bit plane coding only,
+  • the proposed model (spike coding combined with bit plane coding).
+As demonstrated in Table IV, the combination of spike coding with bit-plane coding outperformed other coding methods in terms of validation accuracy.
 As mentioned, using only bit-plane coding resulted in a reduction in accuracy; however, the combination of spike coding with bit plane coding (i.e.
@@
 This could be because among these three datasets, MNIST is the simplest one while Fashion-MNIST is the most complex one.
-Interestingly, the gain of the proposed method compared to the baseline is increased from 0.31% (with MNIST) to roughly 1.4% (with KMNIST and Fashion-MNIST).
+Interestingly, the gain of the proposed method compared to the baseline is increased from 0.31% (with MNIST) to roughly 1.45% (with KMNIST) and averaged at 1.04%.
 That means the impact of bit plane coding is more prominent with more complex datasets.
-Upon further analysis of spike distribution complexities for various coding methods, as presented in Table IV, several insights emerge.
-First, we observe that the baseline method has the lowest complexity in MNIST and KMNIST datasets, while it has the highest complexity in the Fashion-MNIST dataset.
-We suspect that this difference stems from the nature of these datasets.
-Both MNIST and KMNIST primarily use binary pixel values, with zeros representing the background and ones indicating the presence of a character.
-This binary representation results in a simpler spike distribution.
-In contrast, Fashion-MNIST employs a wider range of pixel values to capture the textures of various clothing items, which introduces greater complexity into the spike distribution when using baseline coding methods.
-This observation suggests that more complex visual datasets may lead to even more pronounced differences in spike distribution when using these coding methods.
-Additionally, by introducing a hybrid coding approach that combines baseline and bit plane coding methods, we hypothesize the potential existence of an optimal complexity threshold at which the SNN might achieve peak performance.
-This idea of an optimizable global maximum suggests that complexity thresholds could play a key role in determining SNN effectiveness.
-Based on our data, we propose that measuring spike distribution complexity could be crucial in tuning SNNs to perform optimally across different datasets and tasks.
-V-B2 The effect of bit plane coding on color images
+VI-B2 Comparison of SNN variants
+TABLE V: Comparison of validation accuracy of the baseline and the proposed method on different SNN variants. Performance gain when compared with the baseline is shown in green.
+Accuracy of coding methods per dataset(%)
+MNIST Fashion-MNIST KMNIST CIFAR10 Average
+SNN variant Base-line Proposed method Base-line Proposed method Base-line Proposed method Base-line Proposed method Base-line Proposed method
+SEW-ResNet18 (ADD) 98.62 98.93 (+0.31) 94.13 95.58 (+1.45) 88.83 90.19 (+1.36) 70.69 73.49 (+2.80) 88.07 89.55 (+1.48)
+SEW-ResNet18 (AND) 97.41 98.48 (+1.07) 87.14 89.14 (+2.00) 90.13 94.26 (+4.13) 62.58 70.95 (+8.37) 84.32 88.21 (+3.89)
+SEW-ResNet18 (IAND) 98.39 98.76 (+0.38) 88.92 90.15 (+1.23) 93.23 94.90 (+1.67) 67.28 71.40 (+4.12) 86.96 88.80 (+1.84)
+Spiking-ResNet18 98.25 98.74 (+0.49) 87.47 89.37 (+1.90) 93.23 95.06 (+1.83) 46.15 61.79 (+15.1) 81.28 86.24 (+4.96)
+Spiking-ResNet34 69.02 92.77 (+23.8) 69.59 79.62 (+10.0) 37.45 72.23 (+34.8) 19.28 32.29 (+13.0) 48.84 69.23 (+20.4)
+Spiking-ResNet50 13.45 46.22 (+32.8) 12.80 57.20 (+44.4) 10.44 23.87 (+13.4) 17.94 27.78 (+9.84) 13.66 38.77 (+25.1)
+In addition, we also evaluated our spike coding scheme in different SNN variants.
+We employ only both grayscale and color image datasets in this part, including MNIST, KMNIST, Fashion-MNIST and CIFAR10.
+We benchmarked against multiple variants of Spiking ResNet[26] with LIF surrogate neurons, which have demonstrated competitive performance in prior work, along with AND and IAND variants of SEW-ResNet18[11].
+In shown in Table V, the results reveal that in all of the tested variants, the proposed method consistently achieves a higher accuracy than the baseline.
+The performance gains range from 0.31% (in the case of SEW-ResNet18 (ADD) on MNIST dataset) to 44.40% (in the case of Spiking-ResNet50 on Fashion-MNIST).
+Also, Spiking-ResNet variants tend to underperform when compared with SEW-ResNet18 variants.
+The average accuracy values across all datasets show that SEW-ResNet18 (ADD) is the best variant.
+Fang et al.
+[11] also demonstrated that this variant usually achieves superior performance.
+Therefore, in subsequent experiments, we exclusively utilize this variant for benchmarking.
+VI-B3 The effect of bit plane coding on color images
+TABLE VI: Average validation accuracy on various computer vision dataset of our proposed method.
+Best performances are denoted in bold.
+Performance gain when compare with the baseline is highlighted in green.
 Dataset Accuracy of coding methods (%)
-Baseline Bit plane coding only Proposed method
-CIFAR10[26] 70.69 37.28 73.49
-CIFAR100[26] 38.57 11.28 42.15
-Caltech101[48] 61.67 43.28 64.55
-Caltech256[49] 30.04 21.75 41.51
-EuroSAT[50] 84.39 79.87 88.48
-Imagenette[51] 72.99 48.20 78.68
-Food101[52] 19.62 6.29 36.95
-GTSRB[53] 93.36 95.85 95.35
-TABLE V: Average validation accuracy on various computer vision dataset of our proposed method.
-Dataset Analysis results of coding methods (bits)
-Baseline Bit plane coding only Proposed method
-AKC SE AKC SE AKC SE
-CIFAR10[26] 7530 0.995 6764 0.996 7190 0.995
-CIFAR100[26] 7374 0.995 6710 0.996 7079 0.995
-Caltech101[48] 310k 0.993 260k 0.992 279k 0.992
-Caltech256[49] 307k 0.991 257k 0.990 276k 0.990
-EuroSAT[50] 30.3k 0.953 22.9k 0.944 27.0k 0.949
-Imagenette[51] 334k 0.987 282k 0.994 302k 0.992
-Food101[52] 339k 0.989 288k 0.995 308k 0.993
-GTSRB[53] 308k 0.902 2914 0.223 5791 0.019
-TABLE VI: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE), values in thousands are represented with k (e.g., 22k = 22,000).
-Highest values of AKC are denoted in red and lowest are denoted in blue.
-Dataset Analysis results of coding methods (bits)
-With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
-AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE
-CIFAR10[26] 7190 0.995 7086 0.992 6671 0.972 6634 0.927 6752 0.957 7083 0.986 6857 0.988
-CIFAR100[26] 7079 0.995 6975 0.991 6606 0.974 6551 0.929 6660 0.955 6984 0.987 6769 0.985
-Caltech101[48] 279k 0.992 269k 0.978 231k 0.957 241k 0.905 242k 0.929 268k 0.989 239k 0.971
-Caltech256[49] 276k 0.990 265k 0.976 233k 0.957 243k 0.907 244k 0.933 244k 0.989 242k 0.972
-EuroSAT[50] 27.0k 0.949 26.5k 0.968 25.2k 0.944 25.3k 0.993 25.8k 0.950 26.2k 0.932 25.9k 0.966
-Imagenette[51] 302k 0.992 292k 0.989 267k 0.957 271k 0.901 276k 0.941 288k 0.979 277k 0.976
-Food101[52] 308k 0.993 297k 0.990 273k 0.986 277k 0.908 280k 0.927 294k 0.975 281k 0.969
-GTSRB[53] 5791 0.019 5706 0.006 17.6k 0.489 86.1k 0.462 92.9k 0.499 5704 0.005 5703 0.006
-TABLE VII: Comparison of analysis results with our proposed coding method across all tested color models, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE) measured in bits, values in thousands are represented with k (e.g., 22k = 22,000).
-Highest values of AKC are denoted in red and lowest are denoted in blue.
+Baseline Bit plane coding Proposed method
+CIFAR10[30] 70.69 37.28 73.49 (+2.80)
+CIFAR100[30] 38.57 11.28 42.15 (+3.58)
+Caltech101[55] 61.67 43.28 64.55 (+2.88)
+Caltech256[56] 30.04 21.75 41.51 (+11.47)
+EuroSAT[57] 84.39 79.87 88.48 (+4.09)
+Imagenette[58] 72.99 48.20 78.68 (+5.69)
+Food101[59] 19.62 6.29 36.95 (+17.33)
+Average 53.99 35.42 60.83 (+6.84)
+For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[30], CIFAR100[30], EuroSAT[57], Caltech101[55], Caltech256[56], Imagenette[58] (which is a subset of ImageNet[60]) and Food101[59].
+The default color model of these datasets is RGB.
+Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels.
+So, among these datasets, CIFAR10[30], CIFAR100[30] and EuroSAT[57] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
+The accuracy results of different cases are shown in Table VI.
+We can see that, the accuracy of the proposed method is always higher than that of the baseline.
+The gain ranges from the lowest of 2.80% (with CIFAR100) up to 17.33% (with Food101 dataset), depending on the employed dataset.
+This suggests that the benefit of the proposed method when applying to color images is more significant than when applying to grayscale images.
+TABLE VII: Validation accuracy comparison of proposed method among color models.
 Dataset Accuracy of coding methods (%)
@@
 Food101[59] 36.95 37.71 36.98 38.03 36.41 37.10 35.87
-GTSRB[53] 95.35 51.36 70.68 79.18 80.86 51.78 51.78
-Average 65.15 59.21 62.02 63.36 63.49 59.53 59.61
-TABLE VIII: Comparison of validation accuracy among coding methods.
-For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[26], CIFAR100[26], EuroSAT[50], Caltech101[48], Caltech256[49], Imagenette[51] (which is a subset of ImageNet[54]), GTSRB[53], and Food101[52].
-The default color model of these datasets is RGB.
-Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels.
-So, among these datasets, CIFAR10[26], CIFAR100[26] and EuroSAT[50] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
-Across all tested datasets, our proposed methods consistently outperform the baseline method, with the notable exception of the GTSRB[53] dataset, where bit-plane coding alone appears to yield the best performance without the need for combination with traditional spike coding.
-This observation suggests that while our methods generally enhance performance by improving the representation of spike distributions, certain datasets may favor different coding approaches depending on their visual complexity and specific characteristics.
-In the GTSRB[53] case, it seems that bit-plane coding alone is sufficient to capture the essential features for effective processing.
-The analysis results presented in Table VI highlight the relationship between complexity and performance in different coding methods.
-Bit-plane coding consistently exhibits lower complexities across datasets, while traditional spike coding tends to produce higher complexity values.
-This outcome aligns with our hypothesis that complexity increases in visual datasets with greater detail and variation.
-In further investigating the GTSRB[53] dataset, we observed a significant reduction in spike distribution complexity (over 99% when comparing baseline with proposed scheme and pure bit planes coding in Table VI) alongside improved model performance.
-Usually a decrease in complexity can sometimes indicate a loss of critical information, it may also suggest that the spike distribution has become more streamlined, capturing essential features more effectively.
-However, a complexity reduction at such large scale could also introduce elements of instability (shown in Table VII), as it can be challenging to discern whether the decreased complexity represents a beneficial simplification or a detrimental loss of information.
-Consequently, while our findings indicate promising results, we suggest further exploration of complexity thresholds to ensure consistent performance across varying datasets.
-V-B3 Impacts of color models
-In this part, also we investigate two options, namely 1) the baseline method and 2) proposed method.
-The accuracy results of different cases are shown in Table VIII.
-We can see that, when using RGB color model, the accuracy of the proposed method is always higher than that of the baseline.
-The gain ranges from about 2% (with GTSRB[53] dataset) up to 11% (with Caltech256[49] dataset), depending on the employed dataset.
-This suggests that the benefit of the proposed method when applying to color images is more significant than when applying to grayscale images.
-Interestingly, the results show that RGB is not always the best color model.
-Among the eight datasets, RGB provides the best performance in three cases (namely CIFAR100[26], Caltech256[49], GTSRB[53]).
+Average 60.83 60.43 60.79 61.09 61.01 60.64 60.74
+Figure 4: Validation accuracy visualization of proposed method among color models.
+VI-B4 Impacts of color models
+In this part, we investigate the effect of color models on the performance of proposed method.
+The results are provided in Table VII and also visualized in Fig.
+4.
+In this figure, the baseline model is also included for the sake of comparison.
+From Fig.
+4, it can be seen that our proposed method using different color models is consistently better than the baseline.
+Also, in general, the variations in performances of different color models are small.
+From Table VII, the results show that RGB is not always the best color model.
+Among the eight datasets, RGB provides the best performance in only 2 cases (namely CIFAR100 and Caltech256).
 CMY does not achieve any best results, whereas each of the other five color models achieves the best performance once.
-Especially, with Caltech101[48] dataset, the accuracy of HSV color model is 68.41%, which is much higher than that of RGB color model (64.55%).
-When considering the average of accuracy values across all datasets, the RGB color model is significantly better than other color models (mentioned the last row of Table VIII) and even though there are performance degradation on GTSRB[53], averaged result across all dataset proven to at least out performed baseline method.
-From the above discussion, the key findings of our research can be summarized as follows.
-  • The proposed method provides benefits for both grayscale images and color images.
-  • The gain of the proposed method depends on the employed dataset. For practical color datasets, the gain is from 2% up to 11%.
-  • In terms of average accuracy value over all datasets, RGB is the best color model and CMY is the worst one.
-  • However, considering individual datasets, RGB is not always the best color model. It provides the best performance in only three datasets.
-  • Especially, with the Caltech101 dataset, the performance of HSV color model (68.41%) is much better than that of the RGB color model (64.55%).
-VI Discussion
-VI-A Complexity or entropy?
-In this study, we employ two primary metrics, an approximated Kolmogorov complexity and Shannon entropy, to analyze the effects of bit-plane coding on the distribution of spike trains.
-These metrics were chosen because they are widely used in signal processing and information theory, making them well-suited for analyzing the coded spike and bit-plane signals.
-By treating both coded spikes and bit planes as types of signals, we believe that Kolmogorov complexity and Shannon entropy can provide valuable insights into the complexity of spike distributions.
-While we conducted limited analysis with Shannon entropy, our results indicate a positive correlation between Shannon entropy and Kolmogorov complexity.
-This suggests that Shannon entropy could serve as an alternative to Kolmogorov complexity for those interested in measurement and analysis.
-However, we find Kolmogorov complexity to be more effective, as it provides clearer distinctions between measurements when compared to Shannon entropy.
-Despite achieving substantial insights through these metrics, we recognize that measuring complexity and entropy alone may not fully capture the intricacies of image data.
-Images inherently contain structural complexity and spatial information, which can be difficult to quantify and analyze.
-Focusing solely on spike signals as sequential data may overlook these aspects.
-A helpful analogy is the historical debate in physics regarding the nature of light, with some scientists viewing light as a wave and others as particles.
-It was only with the discovery of wave-particle duality through experiments like the double-slit experiment that a more comprehensive understanding emerged.
-Similarly, viewing spike signals solely as sequential signals could reveal certain patterns but may not capture the full complexity underlying the optimization process in image processing tasks.
-Currently, tools designed specifically for spike signal analysis that incorporate spatial information for image-based tasks are underexplored.
-This gap highlights the need for further research into analysis methods that can integrate both signal and spatial information for more complete insights.
+Especially, with Caltech101 dataset, the accuracy of HSV color model is 68.41%, which is considerably higher than that of RGB color model (64.55%).
+When averaging the accuracy values across all datasets, the HSL color model slightly outperforms the others for both coding method, as indicated in the last row of Table VII.
+The CMY color model, which does not achieve any best results, also has the lowest average performance.
+Further analysis reveals a distinctive characteristic of the HSL model that may contribute to its relatively stable performance.
+HSL encodes the hue component using a 0–359 degree representation, which results in a higher number of spike counts.
+Consequently, the maximum number of bits required for representation increases to $n_{bit}=9$.
+As theoretically formulated in Section V, this greater bit depth might facilitates a higher SNR ratio that enables better performance.
+In contrast, the CMY model operates within a more limited range of $[0,100]$ and thus needs only $n_{bit}=7$ bits to represent the color values.
+Based on the same logic of Section V, this smaller number of bits reduces the gradient magnitude along with signal term for model SNR and thus results in its lower performance.
+To rigorously quantify the computational overhead introduced by our method with different color models, we measure the running times of an image batch on forward propagation for both baseline and the proposed method.
+As shown in Table VIII, the increases in running times depend on datasets and color models, ranging from 20.52% (with CIFAR-10 dataset and CMY color model) to 105.01% (with CIFAR-10 dataset and HSL color model).
+Averaging over all datasets, the increases are from 48.57% (with CMY color model) to 85.35% (with HSL color model).
+Figure 5 shows the relationship between the average gain in accuracy and the average increase in processing time for different color models.
+We can see that any gain in performance requires some increase in processing time.
+Interestingly, while having the smallest performance gain, the CMY color model also has the lowest processing time increase.
+In contrast, the HSL color model has the greatest gain as well as the highest increase in processing time.
+VI-C Discussion
+Based on these observations, the key findings of our analysis are summarized as follows:
+  • The proposed method yields performance improvements for both grayscale and color images by increasing gradient magnitude and improve model SNR during training process.
+  • The gain of the proposed method depends on the employed dataset, ranging from 0.31% to 1.45% in small grayscale image dataset and 2.80% to 17.33% for practical colored image dataset with RGB-based variants.
+  • Proposed method can be generalized for different SNN variants.
+  • Though RGB is the default color model in existing datasets, it is not always the best one, providing highest results in only two datasets.
+  • Based on both theoretical foundation and in terms of average accuracy value over all datasets, HSL is the best color model and CMY is the worst one.
+Yet HSL provides the best performance in only one dataset.
+TABLE VIII: Comparison of model training times between baseline and proposed method using different color models. The values are average training times per epoch.
+Model Variants Processing time by dataset (ms)
+CIFAR-10 CIFAR-100 Caltech101 Caltech256 EuroSAT Imagenette Food101 Average
+Baseline 41.67 43.55 141.25 146.92 66.05 135.71 140.66 102.26
+Proposed (RGB) 63.66 (+52.77%) 65.39 (+50.15%) 235.13 (+66.46%) 240.41 (+63.63%) 116.77 (+76.79%) 225.83 (+67.07%) 234.74 (+66.89%) 168.85 (+65.12%)
+Proposed (CMY) 50.22 (+20.52%) 55.83 (+28.20%) 210.56 (+49.07%) 215.63 (+46.77%) 102.31 (+54.90%) 217.32 (+60.78%) 211.64 (+50.46%) 151.93 (+48.57%)
+Proposed (YCbCr) 61.23 (+46.94%) 62.31 (+43.08%) 234.44 (+65.98%) 242.25 (+64.89%) 118.64 (+79.62%) 226.41 (+67.50%) 229.95 (+63.48%) 167.89 (+64.18%)
+Proposed (HSL) 85.43 (+105.01%) 87.99 (+102.04%) 256.71 (+81.74%) 277.38 (+88.80%) 131.46 (+99.03%) 237.94 (+76.03%) 249.89 (+77.66%) 189.54 (+85.35%)
+Proposed (HSV) 64.36 (+54.45%) 65.37 (+50.10%) 237.17 (+67.91%) 241.10 (+64.10%) 120.21 (+82.00%) 224.03 (+65.74%) 232.85 (+65.54%) 169.30 (+65.56%)
+Proposed (CIE XYZ) 51.12 (+22.68%) 53.45 (+22.73%) 215.11 (+52.29%) 220.92 (+50.37%) 108.71 (+64.59%) 215.65 (+59.54%) 212.06 (+50.76%) 153.86 (+50.46%)
+Proposed (CIE LAB) 50.95 (+22.27%) 52.96 (+21.61%) 213.86 (+51.41%) 219.80 (+49.61%) 105.93 (+60.38%) 217.89 (+61.20%) 215.71 (+53.36%) 153.87 (+50.47%)
+Figure 5: Relationship between average gain in performance (higher is better) and average increase in processing time (lower is better) of different color models.
+Our experimental results demonstrate that bit-plane coding can enhance the performance of SNNs.
+Despite the significant performance gains, the added computational complexity is a limitation of our method.
+As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some applications.
+Also, another limitation is that, similar to recent studies [26, 11], the experiments in our study are still based on software simulations.
+Implementations and experiments on hardware would be definitely needed to understand the actual complexity.
+As our study is the first one to investigate bit plane coding in SNNs, it can be extended in different directions.
+First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs, especially for different machine learning tasks.
+Second, it is obvious that the choice of color model may depend on the task in question, such as medical image analysis or traffic object recognition.
+So, experiments in different application domains to understand the effects of the proposed method and color models for different tasks.
+In addition, the proposed method should be implemented on some hardware platforms and tested in practical scenarios.
 VII Conclusion
-In conclusion, our research introduces a novel coding method for SNN that leverages bit planes derived from various color models of input image data.
-This approach prove to enhances the computational accuracy of SNN without increasing the model size.
-Through extensive experimental validation, we have demonstrated the effectiveness of our coding strategy across multiple computer vision tasks, achieving notable performance gains compared to conventional methods.
-This study is pioneering in its application of color models within the context of spiking neural networks, and we believe that our findings open new avenues for the development of more efficient and effective SNN models.
-Our work sets a foundation for future research and applications, potentially advancing the field of computational neuroscience and artificial intelligence.
-Appendix A CMY or CMYK?
-In the initial stages of our experimental setup and algorithm design, we encountered challenges in selecting the appropriate device-dependent color model for our experiments.
-Specifically, we debated whether to use the CMY or CMYK color model.
-Each has advantages in different applications, but their compatibility with the neural networks used in our research required careful consideration.
-In image processing and printing, the CMYK model was developed to improve print quality by reducing ink usage in darker regions.
-This is achieved by adding a K component, which reduces the CMY values proportionally, leading to enhanced contrast and more economical ink usage in printed images.
-However, when adapting this approach for digital image processing, particularly for direct usage on digital devices, the inclusion of a K channel becomes less beneficial.
-Unlike printed images, digital displays do not require ink density adjustments, making the K channel unnecessary for accuracy.
-For our experiment, this would introduced dimensional conflicts in tested SNN architectures and coding schemes since they are previously optimized for three-channel data inputs (RGB), including a fourth channel would require additional modifications to model architecture, potentially affecting performance and comparability.
-Furthermore, in our image processing pipeline, we observed that converting images to the CMYK color model using standard image processing libraries (e.g., PIL[55]) did not apply undercolor removal during conversion to CMYK model, resulting in a K channel of zero values.
-This posed a unique problem in our training setup: a zeroed K channel in the input data could nullify both the input signal and its gradients when surrogate gradients are used for backpropagation in SNNs.
-Consequently, this would undermine the learning process by reducing the effectiveness of gradient-based updates, complicating model training and ultimately impacting the accuracy of our results.
-Appendix B Tranning progession details
-This section contain the training progression graph for each variant of experiments in Section V-B3 and Section V-B1.
-It is noteworthy from Figure 4 and 5 that performance fluctuations were observed in both traditional coding and our coding method.
-We hypothesize that these fluctuations might related to inherent error rate during sampling process in Poisson-distributed spike trains[3].
-Another hypothesis is that although bit planes may possess a certain degree of spatial information, they can also encompass noise on low order bit planes that interferes with the training process or have little to none distinguishable features (which might explain how on low noises single-channel image dataset such as MNIST, the performance of only using bit planes coding remain competitive to traditional rate coding with less amount of time step).
-Consequently, the selective identification of bit planes that retain valuable spatial information continues to pose a significant challenge.
-(a) MNIST
-(b) Fashion-MNIST
-(c) KMNIST
-Figure 3: Comparison in accuracy progression and loss progression between coding methods on different low-resolution gray-scale image classification datasets.
-(a) CIFAR10[26]
-(b) CIFAR100[26]
-(c) Caltech101[48]
-(d) Caltech256[49]
-(e) EuroSAT[50]
-(f) Imagenette[51]
-(g) Food101[52]
-(h) GTSRB[53]
-Figure 4: Comparison in loss progression of traditional rate coding with our method on image classification datasets.
-On every sub-figure, loss details of every color models is shown on the left and average result of all models on the right with maximum loss difference denoted on top of figure.
-(a) CIFAR10[26]
-(b) CIFAR100[26]
-(c) Caltech101[48]
-(d) Caltech256[49]
-(e) EuroSAT[50]
-(f) Imagenette[51]
-(g) Food101[52]
-(h) GTSRB[53]
-Figure 5: Comparison in accuracy progression of traditional rate coding with our method on image classification datasets.
-On every sub-figure, accuracy details of every color models is shown on the left and average result of all models on the right with maximum accuracy difference denoted on top of figure.
+In this paper, we have introduced a novel coding method for SNN that leverages bit planes of input image data.
+This approach proved to enhance the computational accuracy of SNN without increasing the model size.
+Through providing theoretically justified proof and extensive experimental validation, we have demonstrated the effectiveness of our coding strategy with multiple well known coding scheme across different grayscale and color datasets.
+This study is pioneering in the investigation of bit planes and color models within the context of SNN.
+We expect that the findings in this paper open new avenues for the development of more efficient and effective SNN models.
+In future work, we will further investigate the impacts of bit plane coding and color models in different application domains, such as medical analysis and object detection for traffic smart cameras.
+In addition, the proposed method will be implemented and tested on practical hardware platforms.

### v4

@@
-Improvement of Spiking Neural Network with Bit Planes and Color Models
+
 Abstract
@@
 By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more effective SNNs models in future researches and applications.
-Source code for our experiments is available at https://github.com/luutn2002/bit-plane-snn
 Index Terms:
 Spiking neural network, spike coding, image classification, bit planes, color models
+[…]
+=-15pt
 I Introduction
@@
 $concat()$ Tensor concatenation function, concatenating a list of tensors resulted to a tensor contain spikes of shape ($n_{bit}$, $X_{shape}$)
-$convertColor()$ Tensor color model converting function, convert $X$ to a specified color model decided by value of $colormodel$
-$colormodel$ A string value specify the color model to convert the tensor to (i.e. ”RGB”, ”CMY”)
-$mod$ Modulo division
+$convertColor()$ Tensor color model converting function, convert $X$ to a specified color model decided by value of colormodelcolormodel
+colormodelcolormodel A string value specify the color model to convert the tensor to (i.e. ”RGB”, ”CMY”)
+modmod Modulo division
 $L$ A list-like [16] data structure, used to temporarily store coding result
@@
 Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [21].
-Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [9] or ANN-to-SNN conversion approaches [22].
+Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [9, 11, 22], hybrid SNN training [23] or ANN-to-SNN conversion approaches [24].
 Thus, SNNs present a compelling alternative to ANNs for energy-efficient and biologically plausible computations, while also introducing unique challenges in model design and training.
@@
 Although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still unexplored.
-As far as we concern, our research is first to explore application of bit planes on SNNs spike coding process.
+As far as we are concerned, our research is first to explore application of bit planes on SNNs spike coding process.
 III-B Color models in deep learning
@@
 $X_{max}$ is defined, $X_{max}>0$ and $X\in\mathbb{Z}$
-function BitplaneEncode($X,X_{max}$, $colormodel$)
-     $X=convertColor(X,colormodel)$triangleright Convert $X$ to specified color model
-     $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$
-     $L\leftarrow[\text{ }]$triangleright Empty list initialization
+function BitplaneEncode($X,X_{max}$, colormodelcolormodel)
+  $X=convertColor(X,colormodel)$ $\triangleright$ Convert $X$ to specified color model
+  $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$ $\triangleright$ Get number of bit required to encode image from highest possible value $X$
+  $L\leftarrow[\text{ }]$ $\triangleright$ Empty list initialization
   for each $X_{shape}.C$ do
    for $i\leftarrow 0,i<n_{bit},i++$ do
-              $L.insert(X\mod{2})$triangleright Element-wise modulo of tensor, then store the bit plane to $L$
-              $X\leftarrow\left\lfloor X/2\right\rfloor$triangleright Element-wise division of input image tensor and floor the result
+     $L.insert(X\mod{2})$ $\triangleright$ Element-wise modulo of tensor, then store the bit plane to $L$
+     $X\leftarrow\left\lfloor X/2\right\rfloor$ $\triangleright$ Element-wise division of input image tensor and floor the result
    end for
   end for
-     return $concat(L)$triangleright Concatenate all stored sub tensors in $L$ to a single tensor
+  return $concat(L)$ $\triangleright$ Concatenate all stored sub tensors in $L$ to a single tensor
 end function
@@
 $P_{\mathbf{w}}(\mathbf{x})$ $=\mathbb{E}_{T}\left[p_{\mathbf{w}}(e(\mathbf{x}))\right],$
-where the surrogate gradient model $P_{\mathbf{w}}(\mathbf{x})$ encodes the input $\mathbf{x}$ into a spike signal using an encoding function $e(\mathbf{x})$ and computes the expected output signal across the time dimension $T$ to generate the final stochastic prediction.
+where the surrogate gradient model $P_{\mathbf{w}}(\mathbf{x})$ encodes the input $\mathbf{x}$ into a spike signal using an encoding function $e(\mathbf{x})$ and computes the expected output signal across the time dimension $T$ to generate the final stochastic prediction (basis for this statement can be found ar Appendix A).
 Previous studies in gradient-based optimization[45, 46] have demonstrated a positive correlation between a model’s performance and the signal-to-noise ratio (SNR) of the stochastic gradient, defined by:
@@
 $Noi(\delta(\mathbf{x}))=\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}$
-$\delta(\mathbf{x})=h(\mathbf{x})g_{\mathbf{w}}(\mathbf{x})$
+$\delta(\mathbf{x})=h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x})$
 where $g_{{\mathbf{w}}}({\mathbf{x}})=\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}({\mathbf{x}})$ is the Jacobian of model $p_{{\mathbf{w}}}({\mathbf{x}})$ with respect to model weight $\mathbf{w}$.
 Since this correlation does not depend on input gradient calculation, we can directly apply this in our research.
-To improve the SNR, we can either increase signal term $Sig(\delta(\mathbf{x}))$ or reduce noise term $Noi(\delta(\mathbf{x}))$.
-As shown in [11], the norm $\left\|\delta({\mathbf{x}})\right\|$ is positively correlated with timestep $T$:
-$\left\|\delta({\mathbf{x}})\right\|=\left\|h^{\prime}({\mathbf{x}})\right\|\cdot\left\|\frac{\partial\ell}{\partial O^{l+k-1}}\right\|\cdot\left\|\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)\right\|$ (11)
-$\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)=\sqrt{NT\alpha}$
-$\alpha=\left[\Phi\,\Theta^{\prime}(1-V_{th})^{2k}+(1-\Phi)\,\Theta^{\prime}(0-V_{th})^{2k}\right]$
-$\Phi=\frac{\sum_{j=0}^{N-1}\sum_{t=0}^{T-1}S^{l}_{j}[t]}{N\,T}$
-where the equation denote the magnitude of $\left\|\delta({\mathbf{x}})\right\|$ with target function gradient $h^{\prime}({\mathbf{x}})$, previous blocks’ gradient $\frac{\partial L}{\partial O^{l+k-1}}$ of a $k$ sequential blocks $l$ layers Spiking-ResNet architecture[11] and Heaviside function’s gradient $\Theta^{\prime}$ while transmitting signal $S$ of previous layer output $O$ at firing rate $\Phi$ using $N$ neurons with threshold voltage $V_{th}$.
-Upon performing signal concatenation using bit-plane-based signal $B$ of bit length $N_{bit}$ as in the proposed method, the magnitude of $\left\|\delta({\mathbf{x}})\right\|$ and firing rate $\Phi$ would increase as:
-[…]
-$\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)\Theta^{\prime}\left(B^{l+i}-V_{th}\right)=\sqrt{N(T+N_{bit})\alpha}$
-$\alpha=\left[\Phi\,\Theta^{\prime}(1-V_{th})^{2k}+(1-\Phi)\,\Theta^{\prime}(0-V_{th})^{2k}\right]$
-$\Phi=\frac{\sum_{j=0}^{N-1}\sum_{t=0}^{T-1}S^{l}_{j}[t]+\sum_{j=0}^{N-1}\sum_{t=0}^{N_{bit}-1}B^{l}_{j}[t]}{N\,(T+N_{bit})}$
-and improving the chance of obtaining higher $SNR(\delta(\mathbf{x}))$.
+Upon performing signal concatenation using bit-plane-based signal $B$ of bit length $N_{bit}$ as in the proposed method, given certain condition satisfied, we can have:
+$\mathrm{sup}\left(SNR\left(\delta^{bit}(\mathbf{x})\right)\right)\geq\mathrm{sup}\left(SNR\left(\delta(\mathbf{x})\right)\right),$ (11)
+$\mathrm{sup}\left(SNR(\delta^{bit}(\mathbf{x}))\right)=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta^{bit}(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}},$
+$\mathrm{sup}\left(SNR(\delta(\mathbf{x}))\right)=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}$
+and improving the chance of obtaining higher SNR (more background details and proof for Equation 11 can be found at Appendix B).
 Figure 3: Comparing model gradient’s SNR of the proposed method (red), baseline (blue) and bit plane coding (green) of MNIST dataset across the epochs.
@@
 For the baseline method in the experiments, the Poisson-distributed rate coding [50, 51] is employed with a step size $T=10$ where spike $s_{T,i}$ for i-th pixel at step $T$ with intensity $P_{T,i}$ can be interpreted as:
-$s_{T,i}=\begin{cases}1&\text{if }random(0.0,1.0)\geq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (13)
+$s_{T,i}=\begin{cases}1&\text{if }random(0.0,1.0)\leq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (12)
 To enable efficient training, surrogate gradient methods were applied to all IF neurons.
-We utilized the arctangent surrogate gradient function $S(U)$[11, 10].
+We utilized the arctangent surrogate gradient function S(U)S(U)[11, 10].
 Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[47, 52, 53], computed using mean reduction across mini-batches.
@@
 VI-B2 Comparison of SNN variants
-TABLE V: Comparison of validation accuracy of the baseline and the proposed method on different SNN variants. Performance gain when compared with the baseline is shown in green.
+Table V: Comparison of validation accuracy of the baseline and the proposed method on different SNN variants.
+Performance gain when compared with the baseline is shown in green and best performing model on average is denoted in bold.
 Accuracy of coding methods per dataset(%)
@@
 Yet HSL provides the best performance in only one dataset.
-TABLE VIII: Comparison of model training times between baseline and proposed method using different color models. The values are average training times per epoch.
-Model Variants Processing time by dataset (ms)
-CIFAR-10 CIFAR-100 Caltech101 Caltech256 EuroSAT Imagenette Food101 Average
-Baseline 41.67 43.55 141.25 146.92 66.05 135.71 140.66 102.26
-Proposed (RGB) 63.66 (+52.77%) 65.39 (+50.15%) 235.13 (+66.46%) 240.41 (+63.63%) 116.77 (+76.79%) 225.83 (+67.07%) 234.74 (+66.89%) 168.85 (+65.12%)
-Proposed (CMY) 50.22 (+20.52%) 55.83 (+28.20%) 210.56 (+49.07%) 215.63 (+46.77%) 102.31 (+54.90%) 217.32 (+60.78%) 211.64 (+50.46%) 151.93 (+48.57%)
-Proposed (YCbCr) 61.23 (+46.94%) 62.31 (+43.08%) 234.44 (+65.98%) 242.25 (+64.89%) 118.64 (+79.62%) 226.41 (+67.50%) 229.95 (+63.48%) 167.89 (+64.18%)
-Proposed (HSL) 85.43 (+105.01%) 87.99 (+102.04%) 256.71 (+81.74%) 277.38 (+88.80%) 131.46 (+99.03%) 237.94 (+76.03%) 249.89 (+77.66%) 189.54 (+85.35%)
-Proposed (HSV) 64.36 (+54.45%) 65.37 (+50.10%) 237.17 (+67.91%) 241.10 (+64.10%) 120.21 (+82.00%) 224.03 (+65.74%) 232.85 (+65.54%) 169.30 (+65.56%)
-Proposed (CIE XYZ) 51.12 (+22.68%) 53.45 (+22.73%) 215.11 (+52.29%) 220.92 (+50.37%) 108.71 (+64.59%) 215.65 (+59.54%) 212.06 (+50.76%) 153.86 (+50.46%)
-Proposed (CIE LAB) 50.95 (+22.27%) 52.96 (+21.61%) 213.86 (+51.41%) 219.80 (+49.61%) 105.93 (+60.38%) 217.89 (+61.20%) 215.71 (+53.36%) 153.87 (+50.47%)
+Table VIII: Comparison of model total forward and backward propagation time between baseline and proposed method using different color models.
+The values are average training times per epoch per images batch of 16
+.
+Model Variants Processing time by dataset (ms) CIFAR-10 CIFAR-100 Caltech101 Caltech256 EuroSAT Imagenette Food101 Average Baseline 41.67 43.55 141.25 146.92 66.05 135.71 140.66 102.26 Proposed (RGB) 63.66 (+52.77%) 65.39 (+50.15%) 235.13 (+66.46%) 240.41 (+63.63%) 116.77 (+76.79%) 225.83 (+67.07%) 234.74 (+66.89%) 168.85 (+65.12%) Proposed (CMY) 50.22 (+20.52%) 55.83 (+28.20%) 210.56 (+49.07%) 215.63 (+46.77%) 102.31 (+54.90%) 217.32 (+60.78%) 211.64 (+50.46%) 151.93 (+48.57%) Proposed (YCbCr) 61.23 (+46.94%) 62.31 (+43.08%) 234.44 (+65.98%) 242.25 (+64.89%) 118.64 (+79.62%) 226.41 (+67.50%) 229.95 (+63.48%) 167.89 (+64.18%) Proposed (HSL) 85.43 (+105.01%) 87.99 (+102.04%) 256.71 (+81.74%) 277.38 (+88.80%) 131.46 (+99.03%) 237.94 (+76.03%) 249.89 (+77.66%) 189.54 (+85.35%) Proposed (HSV) 64.36 (+54.45%) 65.37 (+50.10%) 237.17 (+67.91%) 241.10 (+64.10%) 120.21 (+82.00%) 224.03 (+65.74%) 232.85 (+65.54%) 169.30 (+65.56%) Proposed (CIE XYZ) 51.12 (+22.68%) 53.45 (+22.73%) 215.11 (+52.29%) 220.92 (+50.37%) 108.71 (+64.59%) 215.65 (+59.54%) 212.06 (+50.76%) 153.86 (+50.46%) Proposed (CIE LAB) 50.95 (+22.27%) 52.96 (+21.61%) 213.86 (+51.41%) 219.80 (+49.61%) 105.93 (+60.38%) 217.89 (+61.20%) 215.71 (+53.36%) 153.87 (+50.47%)
 Figure 5: Relationship between average gain in performance (higher is better) and average increase in processing time (lower is better) of different color models.
@@
 As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some applications.
+Still, along with most spike coding methods utilized in surrogate training in SNN, both proposed method and baseline are still able to compute spike signal within linear time complexity $\mathcal{O}(n)$[3].
+Empirical evaluation of processing speed and utilized memory could varied across datasets, attributed to factors such as input resolution, hardware specifications, and software library versions.
 Also, another limitation is that, similar to recent studies [28, 11], the experiments in our study are still based on software simulations.
@@
 As our study is the first one to investigate bit plane coding in SNNs, it can be extended in different directions.
-First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs, especially for different machine learning tasks.
-Second, it is obvious that the choice of color model may depend on the task in question, such as medical image analysis or traffic object recognition.
+First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs for different machine learning tasks or cross ANN-SNN benchmarking[63].
+Second, it is obvious that the choice of color model may depend on the task in question, such as temporal-based segmentation[64] , object recognition[65] and knowledge distillation [66].
 So, experiments in different application domains to understand the effects of the proposed method and color models for different tasks.
-In addition, the proposed method should be implemented on some hardware platforms and tested in practical scenarios.
+Additionally for future works, the proposed method should be implemented by mapping proposed spike coding method on hardware platforms for deployment such as FPGA or Intel Loihi chip[6] and tested in practical scenarios such as low power edge devices[67].
 VII Conclusion
@@
 In addition, the proposed method will be implemented and tested on practical hardware platforms.
+Acknowledgments
+We are sincerely thankful for the constructive criticisms and insights of reviewers for the betterment of our research.
+Data availability
+Source code for our experiment is publicly available on GitHub (https://github.com/luutn2002/bit-plane-snn).
+Appendix A Basis for stochastic theorem reusability in SNN surrogate training
+As discussed in Section II-B, surrogate gradient methods attempt to adapt deep learning’s gradient optimization techniques to SNNs, maintaining a close connection to traditional deep learning and convex optimization.
+Given well-established conditions and appropriate modifications, the theoretical principles from these fields can be leveraged to explain the behavior of surrogate gradient methods.
+To prove this viewpoint we consider a classical stochastic optimization problem associated with learning a target function $h$ can be formulated as:
+$\min_{\mathbf{w}}F_{h}(\mathbf{w})=\mathbb{E}_{\mathbf{x}}\left[\ell(p_{\mathbf{w}}(\mathbf{x}),h(\mathbf{x}))\right]$ (13)
+where $\ell$ is a loss function, $\mathbf{x}$ represents stochastic inputs (assumed to be vectors in Euclidean space), and $p_{\mathbf{w}}$ is a differentiable predictor parameterized by $\mathbf{w}$ (e.g., a traditional ANN of a specific architecture).
+It is assumed that $E_{x}[\left\|\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}\right\|^{2}]\leq G(\mathbf{w})^{2}$ for some scalar $G(\mathbf{w})$, and that $F$ is differentiable.
+The variance of the given cost function can be quantified as:
+$Var(\mathcal{H},F,\mathbf{w})=\mathbb{E}_{h}\left\|\nabla F_{h}(\mathbf{w})-\mathbb{E}_{h^{\prime}}\nabla F_{h^{\prime}}(\mathbf{w})\right\|^{2}$ (14)
+Now, we consider an example theorem along with it’s auxiliary lemmas derived from a theoretical analysis of a gradient-based optimization experiment[45]:
+Theorem A.1 (Shalev et al., 2017).
+Let $\mathbf{x}_{1}^{k}$ denote a $k$-tuple $(\mathbf{x}_{1},\ldots,\mathbf{x}_{k})$ of input instances, and assume that each $\mathbf{x}_{l}$ is i.i.d. standard Gaussian in $\mathbb{R}^{d}$. Define
+$h_{\mathbf{u}}(\mathbf{x}_{1}^{k})=\prod_{l=1}^{k}\text{sign}(\mathbf{u}^{\top}\mathbf{x}_{l}),$
+and the objective (w.r.t. some predictor $p_{\mathbf{w}}$ parameterized by $\mathbf{w}$)
+$F(\mathbf{w})=\mathop{\mathbb{E}}\limits_{\mathbf{x}_{1}^{k}}\left[\ell(p_{\mathbf{w}}(\mathbf{x}_{1}^{k}),h_{\mathbf{u}}(\mathbf{x}_{1}^{k}))\right].$
+Where the loss function $\ell$ is either the square loss $\ell(\hat{y},y)=\frac{1}{2}(\hat{y}-y)^{2}$ or a classification loss of the form $\ell(\hat{y},y)=r(\hat{y}\cdot y)$ for some $1$-Lipschitz function $r$.
+Fix some $\mathbf{w}$, and suppose that $p_{\mathbf{w}}(\mathbf{x})$ is differentiable w.r.t.
+$\mathbf{w}$ and satisfies $\mathop{\mathbb{E}}\limits_{\mathbf{x}_{1}^{k}}\left[\|\frac{\partial}{\partial\mathbf{w}}p_{\mathbf{w}}(\mathbf{x}_{1}^{k})\|^{2}\right]\leq G(\mathbf{w})^{2}$.
+Then if $\mathcal{H}=\{h_{\mathbf{u}}:\mathbf{u}\in\mathbb{R}^{d},\|\mathbf{u}\|=1\}$, then
+$\textrm{Var}(\mathcal{H},F,\mathbf{w})\;\leq\;G(\mathbf{w})^{2}\cdot O\left(\sqrt{\frac{k\log(d)}{d}}\right)^{k}.$
+Lemma A.2 (Shalev et al., 2017).
+Let $h_{1},\ldots,h_{n}$ be real-valued functions on some Euclidean space, which belong to some weighted $L_{2}$ space.
+Suppose that $\|h_{i}\|_{L_{2}}=1$ and $\max_{i\neq j}|\langle h_{i},h_{j}\rangle_{L_{2}}|\leq c$.
+Then for any function $g$ on the same domain,
+$\frac{1}{n}\sum_{i=1}^{n}\langle h_{i},g\rangle_{L_{2}}^{2}\,\leq\,\|g\|_{L_{2}}^{2}\left(\frac{1}{n}+c\right).$
+Lemma A.3 (Shalev et al., 2017).
+If $\mathbf{w},\mathbf{v}$ are two unit vectors in $\mathbb{R}^{d}$, and $\mathbf{x}$ is a standard Gaussian random vector, then
+$\left|\mathbb{E}_{\mathbf{x}}\left[\text{sign}(\mathbf{w}^{\top}\mathbf{x})\,\text{sign}(\mathbf{v}^{\top}\mathbf{x})\right]\right|\leq|\langle\mathbf{w},\mathbf{v}\rangle|$
+Theorem A.1 holds true as long as Lemma A.2 and Lemma A.3 remain valid. There are two primary approaches to generalizing this theorem for surrogate gradient methods:
+  • Demonstrating that as the distribution of the input $\mathbf{x}$ have a distribution similar to a Bernoulli distribution, the lemmas remain valid, thereby ensuring the theorem holds.
+  • Reformulating the surrogate gradient problem as a conventional stochastic optimization problem.
+We first consider the first approach.
+Suppose each entry $\mathbf{x_{i}}$ of the vector $\mathbf{x}$ follows a Bernoulli distribution with parameter $p_{i}$, so that the joint probability of $\mathbf{x}$ is given by:
+$Pr(\mathbf{x})=\prod_{i}p_{i}^{x_{i}}(1-p_{i})^{1-x_{i}}.$
+Since the components of Lemma A.2 do not depend on the distribution of $\mathbf{x}$, the lemma remains unchanged.
+However, Lemma A.3 no longer necessarily holds, as any multiplication involving $\mathbf{x}$ can be interpreted as a $d$-dimensional Littlewood–Offord problem[68], which only admits a probabilistic upper bound rather than a deterministic one.
+Consequently, Lemma A.3 becomes unstable and holds only under the additional assumption that the unit vectors also follow a standard Gaussian distribution, i.e., $\mathbf{w},\mathbf{v}\sim\mathcal{N}(0,1)$.
+Under this assumption, $w^{T}x$ and $v^{T}x$ are jointly zero-mean Gaussian with variance 1 and covariance $\mathbb{E}[w^{T}xx^{T}v]=w^{T}v$. Thus,
+$\mathbb{E}\left[\text{sign}(w^{T}x)\text{sign}(v^{T}x)\right]$ $=2\Pr(w^{T}x\geq 0,v^{T}x\geq 0)$ (15)
+$-2\Pr(w^{T}x\geq 0,v^{T}x\leq 0).$
+Using the quadrant probability of bivariate normal distributions, this simplifies to:
+$2\left(\frac{1}{4}+\frac{\sin^{-1}(w^{T}v)}{2\pi}\right)-2\left(\frac{\cos^{-1}(w^{T}v)}{2\pi}\right)$ (16)
+$=\frac{2\sin^{-1}(w^{T}v)}{\pi}.$
+The absolute value of this expression is upper bounded by $|w^{T}v|$, thereby establishing the stability of Lemma A.3 as long as the condition of $\mathbf{w},\mathbf{v}\sim\mathcal{N}(0,1)$ held truth.
+However, proving each theorem along with its auxiliary lemmas under different input distributions shift would be inefficient since the theorem only hold within strictly constructed condition.
+Alternatively, we can approach this by reformulating surrogate gradient method as a traditional stochastic optimization problem. In this case, our objective is defined as:
+$\min_{\mathbf{w}}F_{h}(\mathbf{w})$ $=\mathbb{E}_{\mathbf{x}}\left[\ell(P_{\mathbf{w}}(\mathbf{x}),h(\mathbf{x}))\right],$ (17)
+$P_{\mathbf{w}}(\mathbf{x})$ $=\mathbb{E}_{T}\left[p_{\mathbf{w}}(e(\mathbf{x}))\right],$
+where the surrogate gradient model $P_{\mathbf{w}}(\mathbf{x})$ encodes the input $\mathbf{x}$ into a spike signal using an encoding function $e(\mathbf{x})$ and computes the expected output signal across the time dimension $T$ to generate the final stochastic prediction.
+Using this reformulation, we ensure that as long as the encoding function $e(\mathbf{x})$ is differentiable, all conventional deep learning theorems remain applicable to the surrogate gradient model.
+If $e(\mathbf{x})$ is not differentiable, the theorem holds only if its components and auxiliary lemmas do not require gradient computation with respect to $\mathbf{x}$.
+In similar manner with Equation 10, gradient of coding function $\frac{\partial e(\mathbf{x})}{\partial\mathbf{x}}$ is not required and Equation 10 can be apply to surrogate SNN model as is.
+Appendix B Proof for SNR inequality
+Firstly, we will provide some background knowledge to assist our mathematical derivation of Equation 11.
+Given a SNN model that employ hard-reset IF neurons with backpropagate through time (BPTT)[11, 10, 48, 23, 22], optimization scheme is defined as:
+[…]
+$s_{i}^{(l)}(t)=\Theta\left(u_{i}^{(l)}(t),V_{\text{th}}\right)$
+[…]
+$\frac{\partial\mathcal{L}}{\partial W^{(l)}}=\sum_{t=1}^{T}\frac{\partial\mathcal{L}}{\partial s^{(l)}(t)}\cdot\frac{\partial s^{(l)}(t)}{\partial u^{(l)}(t)}\cdot\frac{\partial u^{(l)}(t)}{\partial W^{(l)}}$
+where:
+  • $u_{i}^{(l)}(t)$ is the membrane potential of neuron $i$ in layer $l$ at time $t\in T$,
+  • $s_{i}^{(l)}(t)$ is the spike output of neuron $i$, similarly with prior layer signal $s_{i}^{(l-1)}(t)$,
+  • $V_{\text{th}}$ is the threshold voltage,
+  • synaptic weight $W^{(l)}$ and bias $b_{i}^{(l)}$,
+  • $\Theta$ is the Heaviside function,
+  • loss function $\mathcal{L}$ and it’s gradient with respect to weight $\frac{\partial\mathcal{L}}{\partial W^{(l)}}$.
+As noted from Appendix A, SNR expression denoted in Equation 10 [45] is adaptable to surrogate SNN model by substitute results from Equation 18 as:
+$SNR(\delta(\mathbf{x}))=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{2}}{\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{2}},$ (19)
+[…]
+where $g_{{\mathbf{w}}}({\mathbf{x}})=\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}({\mathbf{x}})$ is the Jacobian of model $p_{{\mathbf{w}}}({\mathbf{x}})$ with respect to model weight $\mathbf{w}$.
+Under the case of signal concatenation with bit plane coding of length $N_{bit}$, we would have $\delta^{bit}(\mathbf{x})$:
+$\delta^{bit}(\mathbf{x})$ $=h(\mathbf{x})\cdot\Big(\sum_{t=1}^{T}\frac{\partial s^{(l)}(t)}{\partial u^{(l)}(t)}\cdot\frac{\partial u^{(l)}(t)}{\partial W^{(l)}}$ (20)
+$+\sum_{t=1}^{N_{bit}}\frac{\partial s^{(l)}(t)}{\partial u^{(l)}(t)}\cdot\frac{\partial u^{(l)}(t)}{\partial W^{(l)}}\Big)$
+$=h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x})+h(\mathbf{x})\cdot g^{bit}_{\mathbf{w}}(\mathbf{x})$
+$=h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})$
+Equation 11 hold true if the following theorem is true:
+Theorem B.1.
+Assume that bit channels concatenated gradient $g^{bit}_{\mathbf{w}}(\mathbf{x})$ does not explode nor vanish, bounded as:
+$G(\mathbf{w})^{2}>\left\|\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right\|_{\infty}>0$
+for some scalar $G(\mathbf{w})$, where both gradient of baseline SNN $g^{T}_{\mathbf{w}}(\mathbf{x})$ and bit plane code concatenated SNN $g^{bit}_{\mathbf{w}}(\mathbf{x})$ converge toward a reasonable approximation of $h(\mathbf{x})$ and have similar distribution that satisfy:
+$\mathrm{Cov}\left(h(\mathbf{x}),g^{bit}_{\mathbf{w}}(\mathbf{x})\right)=0,\quad\mathrm{sign}(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\right])>0,$
+$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
+there exist:
+$\mathrm{sup}\left(SNR\left(\delta^{bit}(\mathbf{x})\right)\right)\geq\mathrm{sup}\left(SNR\left(\delta(\mathbf{x})\right)\right).$
+where:
+$\mathrm{sup}\left(SNR(\delta^{bit}(\mathbf{x}))\right)=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta^{bit}(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}},$
+$\mathrm{sup}\left(SNR(\delta(\mathbf{x}))\right)=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}$
+Proof.
+In order to prove Theorem B.1, we first need some auxiliary lemmas to be proven.
+Lemma B.2.
+Assuming bit plane concatenated SNN model $p^{bit}_{\mathbf{w}}(\mathbf{x})$ reasonably approximate $h(\mathbf{x})$ such that bit-plane Jacobian $g^{bit}_{\mathbf{w}}(\mathbf{x})$ satisfy:
+$\mathrm{Cov}\left(h(\mathbf{x}),g^{bit}_{\mathbf{w}}(\mathbf{x})\right)=0,\quad\mathrm{sign}(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\right])>0,$
+while bit plane gradient does not explode nor vanish over stochastic estimation, bounded as:
+$G(\mathbf{w})^{2}>\left\|\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right\|_{\infty}>0$ (21)
+for some scalar $G(\mathbf{w})$, there exist:
+$\left\|\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}\geq\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}.$
+Proof.
+In order for the resulted inequality in our lemma exist, we need to prove that under given condition, there exist:
+[…]
+As noted from initial condition in Equation 21, we have:
+$G(\mathbf{w})^{2}>\left\|\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right\|_{\infty}>0$
+$\Leftrightarrow\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\geq 0$
+$\Leftrightarrow\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\right]\right)\cdot\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\geq 0$
+$\Leftrightarrow\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\right]\cdot\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]+Cov\left(h(\mathbf{x}),g^{bit}_{\mathbf{w}}(\mathbf{x})\right)\right)\geq 0$
+$\Leftrightarrow\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\cdot g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\geq 0$
+[…]
+Under the property of expectation’s sum and Minkowski’s sum [69]:
+$\Leftrightarrow\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\geq\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x})\right]\right)$
+$\Leftrightarrow\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right)\geq\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right).$
+We can then use a similar process with infimum inequality of initial condition:
+$G(\mathbf{w})^{2}>\left\|\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right\|_{\infty}>0$
+$\Leftrightarrow\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\leq 0$
+$\Leftrightarrow\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\right]\right)\cdot\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\leq 0$
+$\Leftrightarrow\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\right]\cdot\mathbb{E}_{\mathbf{x}}\left[g^{bit}_{\mathbf{w}}(\mathbf{x})\right]+Cov\left(h(\mathbf{x}),g^{bit}_{\mathbf{w}}(\mathbf{x})\right)\right)\leq 0$
+$\Leftrightarrow\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\cdot g^{bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\leq 0$
+[…]
+$\Leftrightarrow\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\leq\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x})\right]\right)$
+$\Leftrightarrow\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right)\leq\mathrm{inf}\left(\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right)$
+From here we can conclude that there exist:
+$\left\|\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}\geq\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}$
+under given condition. ∎
+Lemma B.3.
+Assuming both gradient of baseline SNN $g_{\mathbf{w}}(\mathbf{x})$ and bit plane code concatenated SNN $g^{bit}_{\mathbf{w}}(\mathbf{x})$ converge toward a reasonable approximation of $h(\mathbf{x})$ and have similar distribution that satisfy:
+$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
+there exist:
+$\mathbb{E}_{\mathbf{x}}\left\|\delta^{bit}(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}=\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}.$
+Proof.
+Since we have the initial condition that:
+$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
+then:
+$g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim g_{\mathbf{w}}(\mathbf{x})$ (22)
+$\Leftrightarrow Var\left(\delta^{bit}(\mathbf{x})\right)=Var\left(\delta(\mathbf{x})\right)$
+$\Leftrightarrow\left\|\delta^{bit}(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}=\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}$
+$\Leftrightarrow\mathbb{E}_{\mathbf{x}}\left\|\delta^{bit}(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}=\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}.$
+From results of Lemma B.2 and B.3, we can conclude that there exist:
+$\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta^{bit}(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta^{bit}(\mathbf{x})\right]\right\|^{2}_{\infty}}\geq\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}$
+$\Leftrightarrow\mathrm{sup}\left(SNR\left(\delta^{bit}(\mathbf{x})\right)\right)\geq\mathrm{sup}\left(SNR\left(\delta(\mathbf{x})\right)\right).$
+On the other hand, there are special situation where the $L_{2}$ norm $\left\|\delta(\mathbf{x})\right\|_{2}$ directly scale with model Jacobian $g_{\mathbf{w}}(\mathbf{x})$ dimensions.
+If target function $h(\mathbf{x})$ is zero mean normally distributed $\mathcal{N}(0,\sigma^{2}_{h})$, $g_{\mathbf{w}}(\mathbf{x})$ have sufficiently small variance and does not saturate, $\delta(\mathbf{x})$ would follow a zero mean normal distribution (denoted $\mathcal{N}(0,\sigma^{2}_{\delta})$) and its norm can be directly decomposed to:
+$\left\|\delta(\mathbf{x})\right\|_{2}=\sqrt{\sum_{i=1}^{n_{1}^{\delta}}\sum_{j=1}^{n_{2}^{\delta}}\delta_{ij}^{2}}.$
+where each squared entry $\delta_{ij}^{2}$ would follow a chi-squared distribution with 1 degree of freedom with $\delta_{ij}^{2}\sim\sigma^{2}_{\delta}\chi^{2}(1)$, the sum of all squared entries follows a chi-squared distribution with $n_{1}^{\delta}n_{2}^{\delta}$ degrees of freedom (by Cochran’s theorem[70]):
+$\sum_{i=1}^{n_{1}^{\delta}}\sum_{j=1}^{n_{2}^{\delta}}g_{ij}^{2}\sim\sigma^{2}_{\delta}\chi^{2}(n_{1}^{\delta}n_{2}^{\delta}).$
+From the expectation property of the chi-squared distribution where $\chi^{2}(n_{1}^{\delta}n_{2}^{\delta})\sim Gamma(\alpha=nk/2,\theta=2/n)$ of $n$ samples with $k$ degrees of freedom, we have:
+$E\left[\sigma^{2}_{\delta}\chi^{2}(n_{1}^{\delta}n_{2}^{\delta})\right]=\sigma^{2}_{\delta}n_{1}^{\delta}n_{2}^{\delta}.$
+Taking the square root gives:
+$\left\|\delta({\mathbf{x}})\right\|_{2}\approx\left|\sigma_{\delta}\right|\sqrt{n_{1}^{\delta}n_{2}^{\delta}}.$
+Thus, $\left\|\delta({\mathbf{x}})\right\|_{2}$ grows asymptotically as $O(\left|\sigma_{\delta}\right|\sqrt{n_{1}^{\delta}n_{2}^{\delta}})$.
+We can also quantify the probability for lower bound and upper bound of $\left\|\delta({\mathbf{x}})\right\|$ using the consequences of the standard Laurent-Massart bounds[71]:
+$k=n_{1}^{\delta}n_{2}^{\delta}$ (23)
+[…]
+$Pr\left(A\right)\geq 1-e^{-k^{\alpha}}\left(\forall\alpha\in\mathbb{R}\right)$
+It is clear that as $k$ increase, lower bound, upper bound of $\left\|\delta({\mathbf{x}})\right\|_{2}$ and probability of $\left\|\delta({\mathbf{x}})\right\|_{2}$ is within the established range also increase.
+[…]

### v5

@@
 I Introduction
-The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[1], largely driven by the developments of deep learning algorithms.
-Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[2].
+The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[wani2020advances], largely driven by the developments of deep learning algorithms.
+Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[nguyen2021review].
 To bridge this gap, spiking neural networks (SNNs) have emerged as a promising alternative, leveraging the temporal dimension of neuronal activity to more closely emulate the behavior of biological brains.
-SNNs utilize discrete spike events to transmit information between neurons, mirroring the asynchronous and event-driven communication observed in biological neurons[3].
-This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[4].
-The inherent sparsity and temporality in coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [5].
+SNNs utilize discrete spike events to transmit information between neurons, mirroring the asynchronous and event-driven communication observed in biological neurons[auge2021survey].
+This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[blouw2019benchmarking].
+The inherent sparsity and temporality in coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [rajendran2019low].
 Despite their potential, SNNs pose significant challenges, particularly in training methodologies and network optimization.
-Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[6].
-Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9, 10, 11], have been developed to address these challenges, enabling the effective training of SNNs.
+Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[nunes2022spiking].
+Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [srinivasan2017spike, liu2021sstdp] and surrogate gradient methods [neftci2019surrogate, eshraghian2023training, fang2021deep], have been developed to address these challenges, enabling the effective training of SNNs.
 Prior research in ANN has explored the application of bit planes (e.g.
-[12]) and color models [13, 14] to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing.
+[vorabbi2023input]) and color models [gowda2019colornet, taipalmaa2020different] to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing.
 Motivated by these advancements, we aim to investigate whether bit planes and color models can similarly improve the performance of SNNs.
@@
 Furthermore, RGB color model, which is the default color model of many datasets, is not always the best one among the evaluated color models in our study.
-A preliminary version of this paper is published in [15], where the benefit of using bit plane coding is proved by three grayscale datasets.
+A preliminary version of this paper is published in [luu2024], where the benefit of using bit plane coding is proved by three grayscale datasets.
 In this extended version, the following new points have been added.
@@
 modmod Modulo division
-$L$ A list-like [16] data structure, used to temporarily store coding result
+$L$ A list-like [abelson1996structure] data structure, used to temporarily store coding result
 $\mathbb{Z}$ The set of all integers
@@
 SNNs can be considered a biologically inspired variant of ANNs that offer a different computational paradigm.
-Unlike traditional ANNs, which process information in a continuous manner using floating-point activations and backpropagation-based learning [17], SNNs rely on discrete events called spikes to encode and transmit information [18].
-This event-driven nature of SNNs allows for energy-efficient computations, making them suitable for neuromorphic hardware implementations [19].
-While the majority of mainstream ANNs architecture operate on floating point data representations, SNNs mimic the sparse, time-dependent processing found in biological neurons [20].
-Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [21].
-Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [9, 11, 22], hybrid SNN training [23] or ANN-to-SNN conversion approaches [24].
+Unlike traditional ANNs, which process information in a continuous manner using floating-point activations and backpropagation-based learning [rumelhart1986learning], SNNs rely on discrete events called spikes to encode and transmit information [maass1997networks].
+This event-driven nature of SNNs allows for energy-efficient computations, making them suitable for neuromorphic hardware implementations [merolla2014million].
+While the majority of mainstream ANNs architecture operate on floating point data representations, SNNs mimic the sparse, time-dependent processing found in biological neurons [gerstner2002spiking].
+Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [lee2018training].
+Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [neftci2019surrogate, fang2021deep, luu2025accuracy], hybrid SNN training [luu2025hybrid] or ANN-to-SNN conversion approaches [rueckauer2017conversion].
 Thus, SNNs present a compelling alternative to ANNs for energy-efficient and biologically plausible computations, while also introducing unique challenges in model design and training.
-One of the earliest neuron models that is still widely applied in modern SNN architectures is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [25].
+One of the earliest neuron models that is still widely applied in modern SNN architectures is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [abbott1999lapicque].
 In this model, the neuron is characterized by its membrane potential $V$:
@@
 In this model, the firing rate increases linearly and indefinitely with the increase in input current.
-While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons[26, 27, 28], challenges still remain in the conversion process.
+While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons[li2021free, deng2021optimal, hu2021spiking], challenges still remain in the conversion process.
 One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches.
-Despite the introduction of both data-driven and model-driven threshold determination methods[29], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts.
+Despite the introduction of both data-driven and model-driven threshold determination methods[diehl2015fast], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts.
 Additionally, IF neurons lack time-dependent memory.
 Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires.
-This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [30].
-Still, recent study have shown that SNN employed with IF neurons are more energy efficient than traditional ANN[31].
-Within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better than leaky integrate-and-fire (LIF) neurons[30], with the SEW-ResNet[11] architecture being one such example.
-Figure 1: Visualization of the extracted bit planes from a sample image in the CIFAR-10 dataset [32].
+This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [zhang2019tdsnn].
+Still, recent study have shown that SNN employed with IF neurons are more energy efficient than traditional ANN[dampfhoffer2022snns].
+Within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better than leaky integrate-and-fire (LIF) neurons[zhang2019tdsnn], with the SEW-ResNet[fang2021deep] architecture being one such example.
+Figure 1: Visualization of the extracted bit planes from a sample image in the CIFAR-10 dataset [alex2009learning].
 Higher-order bit planes (e.g., most significant bits) encode more spatial and structural information, capturing the prominent features of the image.
@@
 II-B Deep Residual Learning in SNNs
-The success of ResNet[33] in deep learning has motivated efforts to extend residual learning to SNN.
-Prior attempts, commonly referred to as Spiking ResNets[28], have largely replicated the architecture of ResNet[33] by replacing ReLU activation functions with spiking neurons.
+The success of ResNet[he2016deep] in deep learning has motivated efforts to extend residual learning to SNN.
+Prior attempts, commonly referred to as Spiking ResNets[hu2021spiking], have largely replicated the architecture of ResNet[he2016deep] by replacing ReLU activation functions with spiking neurons.
 However, this approach presents certain limitations:
   • Inability to achieve identity mapping across neuron models: Residual learning is predicated on the concept of identity mapping, where layers can be trained to approximate an identity function.
-While this is straightforward in ResNets[33] with ReLU activation, it is not always feasible in Spiking ResNets[28].
-Specifically, [11] claimed that for certain spiking neuron models, such as LIF neurons[25] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging.
+While this is straightforward in ResNets[he2016deep] with ReLU activation, it is not always feasible in Spiking ResNets[hu2021spiking].
+Specifically, [fang2021deep] claimed that for certain spiking neuron models, such as LIF neurons[abbott1999lapicque] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging.
 This complicates the realization of identity mapping.
-  • Vanishing/exploding gradients: Even in cases where identity mapping can be achieved, Spiking ResNets[28] are prone to vanishing or exploding gradients during training.
+  • Vanishing/exploding gradients: Even in cases where identity mapping can be achieved, Spiking ResNets[hu2021spiking] are prone to vanishing or exploding gradients during training.
 This occurs because the gradient propagation through multiple layers of spiking neurons may either diminish to zero or grow uncontrollably, complicating the training of deeper networks.
-To address these issues, SEW ResNet[11] was introduced as a replacement design for residual blocks while utilizing IF neurons.
-The SEW block capitalizes on the binary nature of spikes by applying element-wise functions $g$ to the output of the residual mapping $A_{l}[t]$ and the input spike train $S_{l}[t]$[11].
+To address these issues, SEW ResNet[fang2021deep] was introduced as a replacement design for residual blocks while utilizing IF neurons.
+The SEW block capitalizes on the binary nature of spikes by applying element-wise functions $g$ to the output of the residual mapping $A_{l}[t]$ and the input spike train $S_{l}[t]$[fang2021deep].
 This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models.
-In recent researches, using backpropagation methods to train SNNs [34, 10, 9, 11] has become a common practice in neuromorphic computing.
+In recent researches, using backpropagation methods to train SNNs [huh2018gradient, eshraghian2023training, neftci2019surrogate, fang2021deep] has become a common practice in neuromorphic computing.
 In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
@@
 $\frac{\partial H}{\partial\theta}=\frac{\partial\sigma}{\partial\theta}=\frac{e^{\theta-t}}{(e^{\theta-t}+1)^{2}}$
-This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively [11].
-Similar to other new types of neural networks [35, 36, 37], the choice of input data representation may have strong impacts on output performance.
+This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively [fang2021deep].
+Similar to other new types of neural networks [ngo2023survey, nguyen2022evaluation, nguyen2022quantum], the choice of input data representation may have strong impacts on output performance.
 Thus, in this paper we investigate a new way that employs bit plane coding for data representation in SNNs.
@@
 Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance.
-In [12], researchers used bit planes to encode the input layer in Binary Neural Network (BNN).
+In [vorabbi2023input], researchers used bit planes to encode the input layer in Binary Neural Network (BNN).
 They argued that bit planes retain crucial spatial information, enabling a reduction in the number of multiply-accumulate operations (MACs) without compromising the model’s accuracy or robustness.
 This demonstrates the potential of bit planes in streamlining computations while maintaining competitive performance.
-The use of bit plane slicing on RGB images for adversarial robustness was explored in [38].
+The use of bit plane slicing on RGB images for adversarial robustness was explored in [liu2022defending].
 By selecting the most significant bit planes, the authors developed a classification model that was better equipped to handle adversarial attacks.
 This approach capitalizes on the spatial information encoded in different bit planes, which helps the model resist manipulations that would otherwise degrade its performance.
-In the domain of medical imaging, bit plane slicing has also been applied to improve accuracy in breast cancer recognition tasks[39].
+In the domain of medical imaging, bit plane slicing has also been applied to improve accuracy in breast cancer recognition tasks[chen2019breast].
 Researchers found that lower-order bit planes are more susceptible to noise, and by removing them before inputting the data into a convolutional neural network (CNN), they achieved higher classification accuracy.
@@
 In computer graphics applications, understanding and effectively managing color models is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces.
-Color models can be divided into device-oriented models, user-oriented models and device-independent models [40, 41].
+Color models can be divided into device-oriented models, user-oriented models and device-independent models [ibraheem2012understanding, plataniotis2001color].
 Device-oriented color models are influenced by the signals of the devices and the resulting colors are affected by the technology used for display.
@@
 For our study, we mostly investigate cylindrically described models, including HSL and HSV.
-According to [40], these models also inherit other special properties such as device dependence and cylindrical-coordinate representation.
+According to [ibraheem2012understanding], these models also inherit other special properties such as device dependence and cylindrical-coordinate representation.
 Device-independent color models are essential for defining color signals without reliance on the specific characteristics of any particular device or application.
@@
 We will mostly research color models defined by the International Commission on Illumination (CIE), including CIE XYZ and CIE LAB.
-Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks, such as image recognition[42, 13], image segmentation[14] and adversarial attacks prevention[43].
+Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks, such as image recognition[kim2018efficient, gowda2019colornet], image segmentation[taipalmaa2020different] and adversarial attacks prevention[chyad2025exploring].
 Despite extensive researches on the use of color models in ANN, their application in SNN remains unexplored and requires further investigation.
@@
 The granularity of information available in bit planes offers a unique approach to encoding visual data, which could enhance the learning capabilities of SNNs by providing a richer, multi-level representation of the input.
-To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[44], which is a simple dataset containing grayscale images of hand-written numbers.
+To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[deng2012mnist], which is a simple dataset containing grayscale images of hand-written numbers.
 We compared three approaches: traditional rate coding, bit plane coding only, and the combination of both rate coding and bit plane coding.
@@
 The results are provided in Table III.
-Table III: Average validation accuracy comparison on MNIST[44]
+Table III: Average validation accuracy comparison on MNIST[deng2012mnist]
 Coding method Validation Accuracy (%)
@@
 As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding.
-Initially, the process commences with the conventional spike coding procedure applied to the image tensor.
+Initially, the process commences with the conventional spike coding procedure applied to the color converted image tensor (defaulted to RGB if unspecified in our algorithm).
 The bit plane coding algorithm is then employed to transform an input image into binary planes.
@@
 $X_{max}$ is defined, $X_{max}>0$ and $X\in\mathbb{Z}$
-function BitplaneEncode($X,X_{max}$, colormodelcolormodel)
+function BitplaneEncode($X,X_{max}$, $colormodel=\text{"RGB"}$)
   $X=convertColor(X,colormodel)$ $\triangleright$ Convert $X$ to specified color model
@@
 As discussed in Section II-B, surrogate gradient methods attempt to adapt deep learning’s gradient optimization techniques to SNNs, maintaining a close connection to traditional deep learning and convex optimization.
-A classical stochastic optimization problem associated with learning a target function $h$ can be formulated as[45]:
+A classical stochastic optimization problem associated with learning a target function $h$ can be formulated as[shalev2017failures]:
 $\min_{\mathbf{w}}F_{h}(\mathbf{w})=\mathbb{E}_{\mathbf{x}}\left[\ell(p_{\mathbf{w}}(\mathbf{x}),h(\mathbf{x}))\right],$ (8)
 where $\ell$ is a loss function, $\mathbf{x}$ represents stochastic inputs (assumed to be vectors in Euclidean space), and $p_{\mathbf{w}}$ is a differentiable predictor parameterized by $\mathbf{w}$.
-It is assumed that $E_{x}[\left\|\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}\right\|^{2}]\leq G(\mathbf{w})^{2}$ for some scalar $G(\mathbf{w})$, and that $F$ is differentiable[45, 46].
+It is assumed that $E_{x}[\left\|\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}\right\|^{2}]\leq G(\mathbf{w})^{2}$ for some scalar $G(\mathbf{w})$, and that $F$ is differentiable[shalev2017failures, ghadimi2013stochastic].
 By reformulating our surrogate gradient optimization process as a traditional stochastic optimization problem, our objective can be defined as:
@@
 where the surrogate gradient model $P_{\mathbf{w}}(\mathbf{x})$ encodes the input $\mathbf{x}$ into a spike signal using an encoding function $e(\mathbf{x})$ and computes the expected output signal across the time dimension $T$ to generate the final stochastic prediction (basis for this statement can be found ar Appendix A).
-Previous studies in gradient-based optimization[45, 46] have demonstrated a positive correlation between a model’s performance and the signal-to-noise ratio (SNR) of the stochastic gradient, defined by:
+Previous studies in gradient-based optimization[shalev2017failures, ghadimi2013stochastic] have demonstrated a positive correlation between a model’s performance and the signal-to-noise ratio (SNR) of the stochastic gradient, defined by:
 $SNR(\delta(\mathbf{x}))=\frac{Sig(\delta(\mathbf{x}))}{Noi(\delta(\mathbf{x}))},$ (10)
@@
 VI-A Experimental settings
-To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [11].
-The model and experiment process were implemented using PyTorch [47] and SpikingJelly[48].
-Color conversion algorithms are based on implementations from Kornia [49].
-For the baseline method in the experiments, the Poisson-distributed rate coding [50, 51] is employed with a step size $T=10$ where spike $s_{T,i}$ for i-th pixel at step $T$ with intensity $P_{T,i}$ can be interpreted as:
-$s_{T,i}=\begin{cases}1&\text{if }random(0.0,1.0)\leq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (12)
+To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [fang2021deep].
+The model and experiment process were implemented using PyTorch [paszke2019pytorch] and SpikingJelly[doi:10.1126/sciadv.adi1480].
+Color conversion algorithms are based on implementations from Kornia [eriba2019kornia].
+For the baseline method in the experiments, the Poisson-distributed rate coding [kim2018deep, heeger2000poisson] is employed with a step size $T=10$ where spike $s_{T,i}$ for i-th pixel at step $T$ with intensity $P_{T,i}$ can be interpreted as:
+$s_{T,i}=\begin{cases}1&\text{if }\mathcal{N}(0.0,1.0)\leq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (12)
 To enable efficient training, surrogate gradient methods were applied to all IF neurons.
-We utilized the arctangent surrogate gradient function S(U)S(U)[11, 10].
-Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[47, 52, 53], computed using mean reduction across mini-batches.
-Inspired by traditional deep learning approach in model fine-tunning [52, 53], the optimization process is conducted with the Adam algorithm [54], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets.
+We utilized the arctangent surrogate gradient function S(U)S(U)[fang2021deep, eshraghian2023training].
+Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[paszke2019pytorch, radford2021learning, luu2023blind], computed using mean reduction across mini-batches.
+Inspired by traditional deep learning approach in model fine-tunning [radford2021learning, luu2023blind], the optimization process is conducted with the Adam algorithm [kingma2014adam], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets.
 In the first part of our experiments, we only use grayscale image datasets to clearly see the effects of bit plane coding.
 In the second part, popular datasets of color images are used to see the practical effects of both bit plane coding and color models.
-We allocated 80% of each dataset for training and the remaining 20% with deterministic random seeding for validation.
-All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU.
+We allocated 80% of each dataset for training and the remaining 20% with deterministic random seeding [^1] Due to resource constraints, common validate tactics (such as utilizing k-fold for cross validation, etc.) might not be implementable on deep SNN training, where simulation of multiple time step is required and training cost are commonly higher than ANN on GPU devices.
+Enforcing deterministic random seed offer a good trade off for accuracy validation with certain level of bias in training and validating process.
+for validation.
+All datasets were processed with a batch size of 16 images [^2] While our main purpose for using small batch size is due to resource constraints, prior work have shown that model training with small batch size can still yield optimal solution given proper setting implemented[masters2018revisiting].
+for 100 epochs on an NVIDIA RTX 3090 GPU.
 VI-B Experimental Results
@@
 Baseline Bit plane coding Proposed method
-MNIST[44] 98.62 97.87 98.93 (+0.31)
-KMNIST[55] 94.13 92.33 95.58 (+1.45)
-Fashion-MNIST[56] 88.83 86.66 90.19 (+1.36)
+MNIST[deng2012mnist] 98.62 97.87 98.93 (+0.31)
+KMNIST[clanuwat2018deep] 94.13 92.33 95.58 (+1.45)
+Fashion-MNIST[xiao2017fashion] 88.83 86.66 90.19 (+1.36)
 Average 93.86 92.29 94.90 (+1.04)
 First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information.
-For that, we employed three popular datasets of grayscale images, including MNIST[44], KMNIST[55], Fashion-MNIST[56].
+For that, we employed three popular datasets of grayscale images, including MNIST[deng2012mnist], KMNIST[clanuwat2018deep], Fashion-MNIST[xiao2017fashion].
 The results of MNIST dataset were already shown in Table III.
@@
 We employ only both grayscale and color image datasets in this part, including MNIST, KMNIST, Fashion-MNIST and CIFAR10.
-We benchmarked against multiple variants of Spiking ResNet[28] with LIF surrogate neurons, which have demonstrated competitive performance in prior work, along with AND and IAND variants of SEW-ResNet18[11].
+We benchmarked against multiple variants of Spiking ResNet[hu2021spiking] with LIF surrogate neurons, which have demonstrated competitive performance in prior work, along with AND and IAND variants of SEW-ResNet18[fang2021deep].
 In shown in Table V, the results reveal that in all of the tested variants, the proposed method consistently achieves a higher accuracy than the baseline.
 The performance gains range from 0.31% (in the case of SEW-ResNet18 (ADD) on MNIST dataset) to 44.40% (in the case of Spiking-ResNet50 on Fashion-MNIST).
-Also, Spiking-ResNet variants tend to underperform when compared with SEW-ResNet18 variants.
+Also, Spiking-ResNet variants tend to underperform when compared with SEW-ResNet18 variants when benchmarking under spike-based backpropagation (which had been priorly noted by [fang2021deep] due to gradient instability in previous works and Section II-B).
 The average accuracy values across all datasets show that SEW-ResNet18 (ADD) is the best variant.
 Fang et al.
-[11] also demonstrated that this variant usually achieves superior performance.
+[fang2021deep] also demonstrated that this variant usually achieves superior performance.
 Therefore, in subsequent experiments, we exclusively utilize this variant for benchmarking.
@@
 Baseline Bit plane coding Proposed method
-CIFAR10[32] 70.69 37.28 73.49 (+2.80)
-CIFAR100[32] 38.57 11.28 42.15 (+3.58)
-Caltech101[57] 61.67 43.28 64.55 (+2.88)
-Caltech256[58] 30.04 21.75 41.51 (+11.47)
-EuroSAT[59] 84.39 79.87 88.48 (+4.09)
-Imagenette[60] 72.99 48.20 78.68 (+5.69)
-Food101[61] 19.62 6.29 36.95 (+17.33)
+CIFAR10[alex2009learning] 70.69 37.28 73.49 (+2.80)
+CIFAR100[alex2009learning] 38.57 11.28 42.15 (+3.58)
+Caltech101[li_andreeto_ranzato_perona_2022] 61.67 43.28 64.55 (+2.88)
+Caltech256[griffin_holub_perona_2022] 30.04 21.75 41.51 (+11.47)
+EuroSAT[helber2019eurosat] 84.39 79.87 88.48 (+4.09)
+Imagenette[imagenette] 72.99 48.20 78.68 (+5.69)
+Food101[food101] 19.62 6.29 36.95 (+17.33)
 Average 53.99 35.42 60.83 (+6.84)
-For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[32], CIFAR100[32], EuroSAT[59], Caltech101[57], Caltech256[58], Imagenette[60] (which is a subset of ImageNet[62]) and Food101[61].
+For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[alex2009learning], CIFAR100[alex2009learning], EuroSAT[helber2019eurosat], Caltech101[li_andreeto_ranzato_perona_2022], Caltech256[griffin_holub_perona_2022], Imagenette[imagenette] (which is a subset of ImageNet[deng2009imagenet]) and Food101[food101].
 The default color model of these datasets is RGB.
 Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels.
-So, among these datasets, CIFAR10[32], CIFAR100[32] and EuroSAT[59] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
+So, among these datasets, CIFAR10[alex2009learning], CIFAR100[alex2009learning] and EuroSAT[helber2019eurosat] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
 The accuracy results of different cases are shown in Table VI.
@@
 With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
-CIFAR10[32] 73.49 73.73 73.64 73.39 73.06 73.63 74.06
-CIFAR100[32] 42.15 41.31 41.62 41.94 41.91 41.73 42.03
-Caltech101[57] 64.55 65.65 65.53 66.46 68.41 65.30 66.22
-Caltech256[58] 41.51 40.65 40.84 40.55 39.94 40.50 40.79
-EuroSAT[59] 88.48 86.61 88.11 88.74 88.81 88.87 87.48
-Imagenette[60] 78.68 77.35 78.78 78.55 78.50 77.35 78.70
-Food101[61] 36.95 37.71 36.98 38.03 36.41 37.10 35.87
+CIFAR10[alex2009learning] 73.49 73.73 73.64 73.39 73.06 73.63 74.06
+CIFAR100[alex2009learning] 42.15 41.31 41.62 41.94 41.91 41.73 42.03
+Caltech101[li_andreeto_ranzato_perona_2022] 64.55 65.65 65.53 66.46 68.41 65.30 66.22
+Caltech256[griffin_holub_perona_2022] 41.51 40.65 40.84 40.55 39.94 40.50 40.79
+EuroSAT[helber2019eurosat] 88.48 86.61 88.11 88.74 88.81 88.87 87.48
+Imagenette[imagenette] 78.68 77.35 78.78 78.55 78.50 77.35 78.70
+Food101[food101] 36.95 37.71 36.98 38.03 36.41 37.10 35.87
 Average 60.83 60.43 60.79 61.09 61.01 60.64 60.74
@@
 Despite the significant performance gains, the added computational complexity is a limitation of our method.
-As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some applications.
-Still, along with most spike coding methods utilized in surrogate training in SNN, both proposed method and baseline are still able to compute spike signal within linear time complexity $\mathcal{O}(n)$[3].
+As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some edge applications where resource utilization are of higher priority.
+Still, along with most spike coding methods utilized in surrogate training in SNN, both proposed method and baseline are still able to compute spike signal within linear time complexity $\mathcal{O}(n)$[auge2021survey].
 Empirical evaluation of processing speed and utilized memory could varied across datasets, attributed to factors such as input resolution, hardware specifications, and software library versions.
-Also, another limitation is that, similar to recent studies [28, 11], the experiments in our study are still based on software simulations.
+Also, another limitation is that, similar to recent studies [hu2021spiking, fang2021deep], the experiments in our study are still based on software simulations.
 Implementations and experiments on hardware would be definitely needed to understand the actual complexity.
 As our study is the first one to investigate bit plane coding in SNNs, it can be extended in different directions.
-First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs for different machine learning tasks or cross ANN-SNN benchmarking[63].
-Second, it is obvious that the choice of color model may depend on the task in question, such as temporal-based segmentation[64] , object recognition[65] and knowledge distillation [66].
+First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs for different machine learning tasks or cross ANN-SNN benchmarking[deng2020rethinking].
+Second, it is obvious that the choice of color model may depend on the task in question, such as temporal-based segmentation[patel2021spiking] , object recognition[su2023deep] and knowledge distillation [guo2023joint].
 So, experiments in different application domains to understand the effects of the proposed method and color models for different tasks.
-Additionally for future works, the proposed method should be implemented by mapping proposed spike coding method on hardware platforms for deployment such as FPGA or Intel Loihi chip[6] and tested in practical scenarios such as low power edge devices[67].
+Additionally for future works, the proposed method should be implemented by mapping proposed spike coding method on hardware platforms for deployment such as FPGA or Intel Loihi chip[nunes2022spiking] and tested in practical scenarios such as low power edge devices[zheng2021balancing].
 VII Conclusion
@@
 $Var(\mathcal{H},F,\mathbf{w})=\mathbb{E}_{h}\left\|\nabla F_{h}(\mathbf{w})-\mathbb{E}_{h^{\prime}}\nabla F_{h^{\prime}}(\mathbf{w})\right\|^{2}$ (14)
-Now, we consider an example theorem along with it’s auxiliary lemmas derived from a theoretical analysis of a gradient-based optimization experiment[45]:
+Now, we consider an example theorem along with it’s auxiliary lemmas derived from a theoretical analysis of a gradient-based optimization experiment[shalev2017failures]:
 Theorem A.1 (Shalev et al., 2017).
@@
 Since the components of Lemma A.2 do not depend on the distribution of $\mathbf{x}$, the lemma remains unchanged.
-However, Lemma A.3 no longer necessarily holds, as any multiplication involving $\mathbf{x}$ can be interpreted as a $d$-dimensional Littlewood–Offord problem[68], which only admits a probabilistic upper bound rather than a deterministic one.
+However, Lemma A.3 no longer necessarily holds, as any multiplication involving $\mathbf{x}$ can be interpreted as a $d$-dimensional Littlewood–Offord problem[littlewood1939number], which only admits a probabilistic upper bound rather than a deterministic one.
 Consequently, Lemma A.3 becomes unstable and holds only under the additional assumption that the unit vectors also follow a standard Gaussian distribution, i.e., $\mathbf{w},\mathbf{v}\sim\mathcal{N}(0,1)$.
@@
 Firstly, we will provide some background knowledge to assist our mathematical derivation of Equation 11.
-Given a SNN model that employ hard-reset IF neurons with backpropagate through time (BPTT)[11, 10, 48, 23, 22], optimization scheme is defined as:
+Given a SNN model that employ hard-reset IF neurons with backpropagate through time (BPTT)[fang2021deep, eshraghian2023training, doi:10.1126/sciadv.adi1480, luu2025hybrid, luu2025accuracy], optimization scheme is defined as:
 […]
@@
   • loss function $\mathcal{L}$ and it’s gradient with respect to weight $\frac{\partial\mathcal{L}}{\partial W^{(l)}}$.
-As noted from Appendix A, SNR expression denoted in Equation 10 [45] is adaptable to surrogate SNN model by substitute results from Equation 18 as:
+As noted from Appendix A, SNR expression denoted in Equation 10 [shalev2017failures] is adaptable to surrogate SNN model by substitute results from Equation 18 as:
 $SNR(\delta(\mathbf{x}))=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{2}}{\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{2}},$ (19)
@@
 $\mathrm{Cov}\left(h(\mathbf{x}),g^{bit}_{\mathbf{w}}(\mathbf{x})\right)=0,\quad\mathrm{sign}(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\right])>0,$
-$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
+$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x}),$
 there exist:
@@
 […]
-Under the property of expectation’s sum and Minkowski’s sum [69]:
+Under the property of expectation’s sum and Minkowski’s sum [zakon2004mathematical]:
 $\Leftrightarrow\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\right]\right)\geq\mathrm{sup}\left(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x})\right]\right)$
@@
 Assuming both gradient of baseline SNN $g_{\mathbf{w}}(\mathbf{x})$ and bit plane code concatenated SNN $g^{bit}_{\mathbf{w}}(\mathbf{x})$ converge toward a reasonable approximation of $h(\mathbf{x})$ and have similar distribution that satisfy:
-$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
+$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x}),$
 there exist:
@@
 Since we have the initial condition that:
-$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
+$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x}),$
 then:
-$g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim g_{\mathbf{w}}(\mathbf{x})$ (22)
+$g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim g^{T}_{\mathbf{w}}(\mathbf{x})$ (22)
 $\Leftrightarrow Var\left(\delta^{bit}(\mathbf{x})\right)=Var\left(\delta(\mathbf{x})\right)$
@@
 $\left\|\delta(\mathbf{x})\right\|_{2}=\sqrt{\sum_{i=1}^{n_{1}^{\delta}}\sum_{j=1}^{n_{2}^{\delta}}\delta_{ij}^{2}}.$
-where each squared entry $\delta_{ij}^{2}$ would follow a chi-squared distribution with 1 degree of freedom with $\delta_{ij}^{2}\sim\sigma^{2}_{\delta}\chi^{2}(1)$, the sum of all squared entries follows a chi-squared distribution with $n_{1}^{\delta}n_{2}^{\delta}$ degrees of freedom (by Cochran’s theorem[70]):
+where each squared entry $\delta_{ij}^{2}$ would follow a chi-squared distribution with 1 degree of freedom with $\delta_{ij}^{2}\sim\sigma^{2}_{\delta}\chi^{2}(1)$, the sum of all squared entries follows a chi-squared distribution with $n_{1}^{\delta}n_{2}^{\delta}$ degrees of freedom (by Cochran’s theorem[cochran1934distribution]):
 $\sum_{i=1}^{n_{1}^{\delta}}\sum_{j=1}^{n_{2}^{\delta}}g_{ij}^{2}\sim\sigma^{2}_{\delta}\chi^{2}(n_{1}^{\delta}n_{2}^{\delta}).$
@@
 Thus, $\left\|\delta({\mathbf{x}})\right\|_{2}$ grows asymptotically as $O(\left|\sigma_{\delta}\right|\sqrt{n_{1}^{\delta}n_{2}^{\delta}})$.
-We can also quantify the probability for lower bound and upper bound of $\left\|\delta({\mathbf{x}})\right\|$ using the consequences of the standard Laurent-Massart bounds[71]:
+We can also quantify the probability for lower bound and upper bound of $\left\|\delta({\mathbf{x}})\right\|$ using the consequences of the standard Laurent-Massart bounds[laurent2000adaptive]:
 $k=n_{1}^{\delta}n_{2}^{\delta}$ (23)
@@
 It is clear that as $k$ increase, lower bound, upper bound of $\left\|\delta({\mathbf{x}})\right\|_{2}$ and probability of $\left\|\delta({\mathbf{x}})\right\|_{2}$ is within the established range also increase.
+[Uncaptioned image] Luu Trong Nhan (Member, IEEE) received the B.Sc.
+degree and M.Sc.
+degree in Computer Science from The University of Aizu, Japan, in 2024 and 2025 respectively.
+He is currently working as Research Assistant at College of Information and Communication Technology, Can Tho University, Vietnam.
+His research interests include machine learning, artificial intelligent algorithm, computational neuroscience and quantum computing.
+[Uncaptioned image] Luu Trung Duong received the B.E.
+degree in Informatics from Can Tho University (CTU), Vietnam, in 1997 and M.Sc.
+degree in Mobile and Distributed Computer Network from Leeds Beckett University, England, in 2003.
+He has been working at Can Tho University since 1997 to present, starting as a Network Manager and currently is the Director of Center for Digital Transformation and Communication.
+His research interests include large scale networking solution design, computer security and network management.
+[Uncaptioned image] Pham Ngoc Nam (Member, IEEE) received the bachelor’s degree in electronics engineering from the Hanoi University of Science and Technology (HUST), in 1997, and the M.S.
+degree in artificial intelligence and the Ph.D.
+degree in electrical engineering from KU Leuven, Belgium, in 1999 and 2004, respectively.
+He is currently the Vice Dean of the College of Engineering and Computer Science, VinUniversity, Vingroup.
+He is also a Visiting Scholar with Cornell University.
+He has been the PI of one key national project and three ministerial-level projects.
+He is the author or coauthor of 100 scientific articles, including more than 30 ISI and Scopus publications.
+His research interests include artificial intelligence, QoS/QoE management for multimedia applications, reconfigurable computing, and low-power embedded system design.
+He has been a key member of four other national projects.
+[Uncaptioned image] Truong Cong Thang (Senior Member, IEEE) received the B.E.
+degree from Hanoi University of Technology, Vietnam, in 1997 and Ph.D.
+degree from KAIST, Korea, in 2006.
+From 1997 to 2000, he worked as an engineer in Vietnam Post & Telecom (VNPT).
+From 2006 to 2011, he was a Member of Research Staff at Electronics and Telecommunications Research Institute (ETRI), Korea.
+He was also an active member of Korean delegation to standard meetings of ISO/IEC and ITU-T from 2002 to 2011.
+Since 2011, he has been an Associate Professor of University of Aizu, Japan.
+His research interests include multimedia networking, image/video processing, content adaptation, IPTV and MPEG/ITU standards.
 […]